### v1

# definition definition

[…]
definition definition
[…]
Mendelson and Zhu
[…]
Online Lending Information Acquisition Strategies
[…]
Optimal Information Acquisition Strategies: The Case of Online Lending
[…]
We first examine this question for an online lender that issues unsecured loans to maximize its expected NPV by dynamically adjusting the amounts it lends at an exogenous interest rate. We find that for this model, a “lean experimentation” strategy is optimal. However, when the interest rate is endogenously determined by the lender, either approach may be optimal depending on the monotonicity properties of the demand elasticity. When the demand elasticity is increasing or constant, the lender performs a single “grand experiment” and then continues to offers the same loan amount and interest rate in each period. We study consumer segmentation and quantify the performance of the different regimes. We show that higher income variability can benefit the lender by enabling more effective experimentation.
[…]
1 Introduction and Literature Review
Information Technology (IT) is playing a key role in the transformation of financial services, as is widely recognized by practitioners and information systems (IS) researchers (Xue et al. (2011), Heidemann et al. (2013), Hendershott et al. (2017), Ho et al. (2017), Burtch et al. (2018), Milian et al. (2019), Jiang et al. (2020), Fu et al. (2021), Yang et al. (2022), Deng et al. (2023), JP Morgan Chase & Co. (2023)). Within this broader transformation, online lending has been a particular focus of study in IS and beyond. Online lending received significant attention for several reasons. Lending is inherently an information activity, as lenders have to predict repayment probabilities using diverse data sources (Yum et al. (2012), Cai et al. (2016)). Developments in information and communication technologies (ICT) provide new sources of information and the technology to analyze them, making the lending market a leading user of data science and AI (Lin et al. (2013), Liu et al. (2015), Hu et al. (2023)). Online lending platforms have introduced new business models that take advantage of these developments and in turn use innovative information architectures that significantly differ from those used in conventional lending environments (Ge et al. (2021), Liu et al. (2023b)). The consequent growth of online lending is well-documented (e.g., Buchak et al. (2016), Buchak et al. (2018)) and won’t be reviewed here.
In its comprehensive review of the online lending marketplace, the U.S. Treasury identified two core operating models (Buchak et al. (2016), p.6): the Direct Lender (DL) model, where lenders “hold loans on their balance sheets” and “generate the majority of their revenue through interest income and fees earned on loans”, and the Platform Lender model, where the online lender operates a marketplace that matches lenders to borrowers and facilitates transactions among them. Under both models, loans are typically unsecured (Buchak et al. (2016)). DLs are also known as balance sheet lenders and platform lenders are best known as Peer-to-Peer (P2P) lenders. Both models are stylized representations of the actual operations of most online lenders, which include additional features such as special types of loans, loan securitization and mixed models. However, the two “pure” models provide useful benchmarks for both analytic and policy purposes.
Online lending became popular in the U.S. through the introduction of P2P lending by Prosper (Mendelson (2006)) and Lending Club (LC), grew quickly in China and Europe and attracted much of the attention of IS researchers. Zhao et al. (2017) and Basha et al. (2021) offer comprehensive surveys of the P2P lending market, and we’ll mention here just a few of these studies. Wei and Lin (2017) compare different mechanisms including fixed price and auction for pricing P2P loans, finding that the fixed price mechanism increases the probability of funding. Lin et al. (2017) explore the determinants of default risk in P2P lending. Jiang et al. (2018) study herding behavior on P2P platforms, showing it is influenced by platform market share and funding amounts. Liu et al. (2020) build a theoretical model to study the impact of soft information in P2P markets. Wu et al. (2021) show how competition and uncertainty can induce biased information when platforms report market conditions to investors. Bao et al. (2023) use a socio-technical framework to evaluate risks and propose mitigation strategies for P2P platforms. Wang et al. (2023) examine how lender competition affects the disclosure of approval odds to borrowers. Gopal et al. (2024) use data from LC to develop strategies that make online lending platforms more appealing to institutional investors and Gopal et al. (2023) focus on the inclusiveness of these platforms.
[…]
Another aspect of LC’s transformation was becoming a bank. As a bank, LC takes deposits and serves the broader financial lives of its customers, which allows it to better assess their creditworthiness, as well as lower its DL funding cost. As a result, when LC underwrites loans it has both traditional sources of banking information and the information available to online lenders, making it what we call a hybrid lender. In addition to studying DL, this paper studies and quantifies the value of the hybrid business model.
[…]
As the case of WeBank illustrates, the path to a hybrid business model did not always originate from P2P lending. 007fenqi, an online DL focused on Chinese students, offered deposit and investment services as well as electronic commerce and other services from its 2015 inception (Leong et al. (2017)). In the US, Upstart was founded in 2012 to become the leading AI-based DL. Upstart used its own balance sheet to fund loans, employing AI to assess creditworthiness. While Upstart continues to operate as a DL, it started offering its technology and AI underwriting engine to credit unions and banks, turning them into hybrid lenders. By 2024, Upstart offered these services to more than 100 partners (Upstart (2024)).
In this paper, we model and solve for the optimal information acquisition strategy of an online DL, considering both exogenous and endogenous interest rates. We characterize the solution and analyze its implications. The key determinant of the optimal policy structure is the monotonicity of the demand elasticity with respect to the interest rate. When demand is completely inelastic (a fixed interest rate), the optimal information acquisition policy is characterized by a large number of small, progressive experiments. We call this strategy a Lean Experimentation (LE) strategy. Conversely, when the demand elasticity is constant or decreases with the interest rate, the optimal information acquisition policy a two-step strategy whereby the lender conducts a single grand experiment (GE); then, if the borrower repays, the same loan amount and rate are applied in all subsequent periods without further experimentation. The information acquisition structure also shapes consumer segmentation into unprofitable, profitable, and creditworthy categories. Under LE, the lender is able to profit even from some consumers who ultimately default whereas under GE, only creditworthy consumers are profitable. Counter-intuitively, our findings show that greater income variability can increase profitability by enabling more effective experimentation.
We further introduce a hybrid information architecture that combines dynamic and traditional forms of information acquisition. This model is more profitable than pure DL. Further, under the exogenous interest rate scenario, the advantages of the hybrid architecture over DL intensify as the interest rate decreases: the hybrid model performs better under less profitable market conditions. When the DL can dynamically adjust the interest rate, the advantages of the hybrid architecture over DL increase with the demand elasticity. Our comparison of dynamic and hybrid information architectures shows that the latter not only increases lender profitability but (in the constant elasticity case) also enhances financial inclusion, demonstrating that the two are not mutually exclusive.
Another objective of this paper, which transcends our online lending context, is to analyze how decision-makers facing sequential opportunities to acquire new information in an uncertain environment should do that. Conventional wisdom suggests that they should take advantage of sequential opportunities by conducting multiple small experiments, each building incrementally on the results of earlier ones. Entrepreneurs and product developers are urged to engage in such a “lean startup” (Reis (2011), Blank (2013)) or “lean experimentation” (LE) approach. Furr and Furr (Furr (2022), Chapter 23) argue for making multiple small experiments (10,000 in the case of National Geographic), citing Match.com, Thomas Edison and Amazon.com as masters of LE. Thomke (2003) shows that product development benefits from a process which is based on many small experiments due to the declining cost of experiments. Citing examples from Booking.com, Expedia, Microsoft and Netflix, Thomke (2020) argues that companies should build a culture that enables many small experiments: “It’s actually less risky to run a large number of experiments than a small number.” Moogk (2012) and Giardino et al. (2014) propose that by building a series of minimum viable products, essentially small product prototypes, startups can receive fast market feedback and evolve quickly. Kerr et al. (2014) and Nanda and Rhodes-Kropf (2016) recommend that venture capital firms run a large number of initial experiments by funding multiple smaller-sized startups. Thomke and Bell (2001) address the product development, testing and rework problem as a tradeoff. Among other results, they derive an interesting EOQ-like formula whereby testing frequency is based on the tradeoff between avoidable costs and the fixed cost of testing.
Online lending is an attractive context for studying this question because of its dynamic and information-rich nature. As discussed earlier, the LE strategy is optimal under inelastic demand. However, when the interest rate is endogenously determined and the demand elasticity is increasing or constant, the opposite is true and a single grand experiment (GE) is optimal. This failure of the common LE wisdom calls for a more nuanced analysis of information acquisition strategies.
As most modern innovations entail the development of IS, the choice of GE vs. LE is becoming increasingly important in the IS development area. Modern software development methodologies range between continuous integration (CI) and continuous Deployment (CD) (for example, Casperson (2020)). Whereas CI focuses on testing the integration of code changes within the software development process, CD automates the entire testing and deployment process all the way to live production, enabling systems to quickly adapt to changes in the external environment. Thus, it may be considered a flavor of LE or a facilitating technology, depending on the context.
In what follows, Section 2 introduces our base model featuring an exogenous interest rate and derives the solution and its implications. This Section also presents the optimal strategy under the hybrid architecture and compares its profitability with that of the dynamic model. Section 3 extends the two models (dynamic and hybrid) to endogenous interest rates, solves them and studies their implications. Section 4 examines the impact of income variability on profitability and Section 5 offers our concluding remarks.
2 Base-Case Model
2.1 Model and Results
We model the DL’s problem as a discrete-time, infinite-horizon, dynamic control problem with the following timeline. A prospective borrower (consumer) is characterized by her periodic disposable income (or ability to pay), $\theta$, drawn from a known distribution ${F}$. We call $\theta$ “income.” The lender extends an unsecured loan, as is typically the case in DL, with a periodic interest rate $r$ which is set by the market (or, as is often the case, by regulation; in Section 3 we extend the model to allow for an endogenous interest rate). The corresponding discount rate is $d=\frac{1}{1+r}$. The realization of $\theta$ is unknown to the lender, who knows only the income distribution $F$. We assume $F$ has a differentiable density $f$ defined over $[0,u)$ where $u>0$ may be finite or infinite (our model extends to distributions with positive support $[l,u)$ for $l>0$). The hazard function corresponding to the income distribution ${F}$ is $H(x)=\dfrac{f(x)}{1-F(x)}$. We assume:
Assumption 2.1
$G(x)=xH(x)$ is strictly increasing for all $x\in[0,u)$, $G(0)=0$ and $G(u)\geq 1$.
(When $u$ is infinite, we interpret the condition $G(u)\geq 1$ as $\lim_{x\to\infty}G(x)\geq 1$). Assumption 2.1 holds for many distributions of interest. It clearly holds for a distribution $F$ with a (weakly) increasing hazard rate with finite support $[0,u)$, $f(0)<\infty$ and $\lim_{x\to u}f(u)>0$, such as the uniform distribution. Other distributions satisfying Assumption 2.1 include the Beta, Gamma and Weibull distributions (see in Proofs Section A.1). Not all distributions satisfy Assumption 2.1. For example, the folded Cauchy distribution has a non-monotone $G(x)$.
The consumer applies for a loan at ${t=0}$ and the lender then decides on the initial loan amount ${\mu_{0}}$. The consumer receives income $\theta$ at the end of each period including the end of period 00 (which is also the beginning of period $1$) and pays back the loan amount plus interest, ${y_{0}=(1+r)\mu_{0}=\mu_{0}/d}$, if ${\mu_{0}/d\leq\theta}$. If ${\mu_{0}/d>\theta}$, the consumer defaults and leaves the system. This process continues in the subsequent periods: at the beginning of each period $t$, the lender decides on the loan amount $\mu_{t}$ and at the end of the period, the consumer repays ${\mu_{t}/d}$ if ${\mu_{t}/d\leq\theta}$ or defaults if ${\mu_{t}/d>\theta}$. Figure 1 shows the timeline.
[…]
Figure 1: Model Timeline.
While our model is highly stylized, its assumptions capture key features of the DL market: the lender gradually learns what the consumer can repay and dynamically adjusts the loan amounts in each period, building longer-term relationships with borrowers it deems profitable. This setup is a canonical example of costly information acquisition in a risky environment, where the lender risks losing the entire loan amount in the hope of discovering profitable customers.
[…]
where we define $\Pi^{-1}_{s=0}\operatorname{\mathbb{P}}(\theta\geq y_{s}|\mathcal{H}_{s})=1$.
2.1.1 Structure of the Optimal Policy
If the consumer has not defaulted by time $t$, the conditional distribution of $\theta$ given history $\mathcal{H}_{t}$ is $F(\theta|\theta\geq\max_{-1\leq i<t}\ y_{i})$, with $y_{-1}$ defined as 00. We define the system state as $x_{t}=\max_{-1\leq i<t}\ y_{i}$, the period-$t$ lower-bound on the consumer’s income. The lender’s decision problem is thus Markovian with decisions depending on $x_{t}$. The Bellman equation for our dynamic control problem is
${J(x_{t})\ =\ \max_{y_{t}}\{\ -dy_{t}+\rho\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})(J(x_{t+1})+y_{t})\},}$ (2)
which satisfies the following Lemma.
Lemma 2.2
If an optimal policy $\{y_{t}\}$ exists, it is weakly increasing, i.e., $y_{t}\geq y_{t-1}$ for all $t>0$.
Intuitively, when the lender knows that $\theta\geq y_{t-1}$, setting $y_{t}<y_{t-1}$ is sub-optimal as it could increase the loan amount to $\mu_{t}=\mu_{t-1}=dy_{t-1}$ knowing that the consumer would pay back $y_{t}$, which strictly dominates setting $y_{t}<y_{t-1}$. Further, when the support of $F$ is continuous, Lemma 2.2 implies that if the consumer does not default by the beginning of period ${t}$, $x_{t}=\max_{-1\leq i<t}y_{i}=y_{t-1}$ and the lender updates the lower bound on income to $y_{t}$. Since ${F}$ has a continuous positive support on $[0,u)$, it follows that
${J(x_{t})\ =\ \max_{y_{t}\in[x_{t},u)}\ \ \{x_{t}-dy_{t}+\rho\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})J(y_{t})\}\ =\ \max_{y_{t}\in[x_{t},u)}\ \ {x_{t}-dy_{t}+\rho\dfrac{1-F(y_{t})}{1-F(x_{t})}J(y_{t})}.}$ (3)
As we show in Theorem 2.4, an optimal policy exists with the lender increasing the repayment amounts towards an upper limit $\bar{x}$ defined below.
Lemma 2.3
There exists a unique finite solution $0<\bar{x}<u$ to the equation:
${G(\bar{x})=-\frac{(1-\rho)d}{(1-d)\rho}+1.}$ (4)
The solution $\bar{x}$ increases in $\rho$ and decreases in $d$.
We are now ready to state the structure of the optimal policy.
Theorem 2.4
Optimal policy: there exists a unique optimal policy whereby the lender increases the repayment amounts $y_{t}$ in each period $t$ as long as the consumer repays her loans. The sequence of repayment amounts $\{y_{t},t=0,1,2,\ldots\}$ is strictly increasing to the limit ${\bar{x}}$.
While we are not aware of empirical work in the DL context we study, the results Theorem 2.4 are consistent with empirical findings on increasing loan amounts in commercial lending (Petersen and Rajan (1994), Angelini et al. (1998), Cenni et al. (2015)).
Under Theorem 2.4 the lender follows a classic Lean Experimentation (LE) strategy: it keeps acquiring information by gradually increasing the loan (and corresponding repayment) amounts, with the increases becoming infinitesimally small. The lender’s LE strategy follows a particular structure driven by the threshold $\bar{x}$, which is the indifference point between ceasing experimentation and acquiring new incremental information. To see why, consider two alternative strategies when the lender is in state $x$: (i) play it safe and just exploit the information embedded in $x$ by lending $dx$ (with repayment amount $x$) in each period, resulting in a zero probability of default; vs. (ii) acquire additional information by increasing the repayment amount to $x+\Delta x$ (lending $d(x+\Delta x)$) to further learn about $\theta$; then, if the consumer repays the loan, the lender keeps setting the repayment amounts $y_{t}$ at $x+\Delta x$ in all subsequent periods. The lender’s expected going forward NPV under strategy (i) is
${-dx+\dfrac{\rho(1-d)}{(1-\rho)}x,}$ (5)
where the first term is the current loan amount and the second is the present value of all future interest cash flows. The expected NPV for strategy (ii) to a first order in $\Delta x$ is given by
${-d(x+\Delta x)+\left(1-H(x)\Delta x\right)\dfrac{\rho(1-d)}{(1-\rho)}(x+\Delta x),}$ (6)
where the first term is the current loan amount, $(1-H(x)\Delta x)$ is the probability that the consumer repays in the next period, and $\frac{\rho(1-d)}{(1-\rho)}(x+\Delta x)$ is the present value of the future interest cash flows conditional on no default. The lender’s first-order incremental NPV from following strategy (ii) over (i) is thus
${\frac{\rho(1-d)}{(1-\rho)}\left(1-\frac{(1-\rho)d}{(1-d)\rho}-G(x)\right)\Delta x,}$ (7)
and $\bar{x}$ is the point of indifference between (i) and (ii).
We call a consumer creditworthy if she will never default under the optimal policy, namely if $\theta\geq\bar{x}$. By Theorem 2.4, the optimal policy is extremely conservative: even when the lender encounters a creditworthy consumer, it never lends more than $d\bar{x}$. As $y_{t}$ approaches $\bar{x}$, the lender knows that $\theta\geq\bar{x}$ (i.e., that the consumer will repay) almost surely. Nevertheless, it never sets the repayment amount above $\bar{x}$. Why does this happen? The reason lies in the difference between statistical learning, which focuses on estimating the value of $\theta$, and optimizing net present value, where the the cost of information acquisition – the losses the lender may incur to learn about $\theta$ – are all important. For statistical learning, binary search would lead to the fastest convergence to the true value of $\theta$. However, such learning is costly as the lender has to balance the potential loss of the entire principal amount when it lends more against the smaller, interest-only potential gain. The lender’s loss function is thus highly asymmetric, which results in a conservative information acquisition policy.
2.2 Implications
2.2.1 Lean Experimentation Strategy
Our results lend support to the Lean Experimentation (LE) strategy structure: the lender keeps experimenting by gradually increasing its loan amounts to identify how much more it can lend. This is best shown by considering the case of uniform income distribution, $F\sim U[0,1]$, where we have a closed-form solution:
Theorem 2.5
When $F\sim U[0,1]$, there is a unique solution to the Bellman equation (3) with lending threshold
${\bar{x}=\dfrac{(\rho-d)}{2\rho-d-d\rho}\in(0,1).}$ (8)
The optimal repayment amounts $y_{t}$ increase linearly in the state $x$ to the limit $\bar{x}$ with $y_{t}(x_{t})=mx_{t}+n$, where
$m$ $=\dfrac{d}{2\rho(1-a)},\quad n=\dfrac{(\rho-d+b\rho)}{2\rho(1-a)},\textup{ and}$ (9)
$a=$ $\dfrac{1}{2}\left(1-d\sqrt{\dfrac{\rho-d^{2}}{\rho d^{2}}}\right),\quad b=\dfrac{1}{2\rho-\rho d-d}\left((\rho-d)\left(d\sqrt{\dfrac{\rho-d^{2}}{\rho d^{2}}}+d-1\right)\right),$
$c=$ $\dfrac{1}{2(1-\rho)(2\rho-d-\rho d)^{2}}\left((\rho-d)^{2}\left(1-2d+\rho-(1-\rho)\sqrt{1-\frac{d^{2}}{\rho}}\right)\right).$
Figure 2 shows that the optimal policy is a classic case of LE: in each state $x_{t}$, the lender slightly increases the loan amount to test whether the consumer can repay $y_{t}(x_{t})>x_{t}$. The amounts at risk, $(y_{t}(x_{t})-x)$, keep decreasing. They are always positive but tend to zero at the limit.
Figure 2: Optimal lending policy $\{y^{*}_{0},y^{*}_{1},y^{*}_{2},\dots\}$ for uniform income distribution with $\rho=0.95$ and $d=0.833\ (r=20\%)$. The x-axis shows the current state (lower bound on income). On the y-axis, $y_{t}(x_{T})$ is the optimal repayment amount when the current state is $x$.
The results in this Section provide a theoretical foundation for the “lean startup” or LE approach advocated by practitioners. An alternative foundation for LE may be found in the multi-armed bandit (MAB) model (Katehakis and Veinott Jr (1987), Lai et al. (1985)). In this model, each action (arm) is associated with a random reward following a probability distribution which is unknown a-priori, and in each period the decision-maker needs to choose an action so as to maximize the expected total payoff over the decision horizon. Our model cannot be analyzed using the MAB approach as the lending market is more complex than the MAB environment. (Specifically, the MAB model requires that the action selected today won’t influence the returns on any of the arms in future periods, whereas in the lending context the actions taken today greatly influence future returns). However, it is worth pointing out that the MAB model can provide a foundation for LE strategies in other models, a fact which to our knowledge has not been recognized in the “lean startup” or experimentation literature. Under the MAB model, the optimal policy gradually increases the emphasis on immediate returns (“exploitation”) at the expense of experimentation. Within that framework, the UCB algorithm (Agrawal (1995), Lattimore and Szepesvári (2020)) which is asymptotically optimal for the MAB problem, gives rise to an LE-like structure.
2.2.2 Consumer Segmentation under Lean Experimentation
[…]
The NPV of a type-$k$ consumer comprises (i) the NPV of her interest payments in periods $0,1,2,\ldots,k$; and (ii) the (negative) present value of the final default loss in period $k$ (zero when $k=\infty$). Component (i) is given by ${PV^{R}_{k}=\sum^{k-1}_{i=0}\rho^{i}(\rho-d)y_{i},}$ and (ii) is ${PV^{I}_{k}=\rho^{k}dy_{k},}$ so the overall NPV of a type-$k$ consumer is ${\Pi_{k}=PV^{R}_{k}-PV^{I}_{k}=\sum^{k-1}_{i=0}\rho^{i}(\rho-d)y_{i}-\rho^{k}dy_{k}.}$ For a creditworthy consumer ($k=\infty$), this reduces to $\Pi_{\infty}=\sum^{\infty}_{i=0}\rho^{i}(\rho-d)y_{i}$.
Importantly, the lender can make money on a non-creditworthy consumer when the present value of her interest payments exceeds that of the default loss. We thus split the non-creditworthy consumers into profitable and unprofitable based on their NPV, which results in three categories:
  • Unprofitable consumers: all type-$k$ consumers with $\Pi_{k}<0$,
  • Profitable consumers: all type-$k$ consumers (for a finite $k$) with $\Pi_{k}\geq 0$, and
  • Creditworthy (type-$\infty$) consumers.
Figure 3 shows how consumers’ NPVs (red lines) vary with their types $k$ for different interest rates, with the NPVs decomposed into those of the interest payments (in blue) minus the default loss (shaded). With a high interest rate (lower $d$, Figure 3(a)), the NPV increases with the consumer’s income, as might be expected: higher-income consumers generate more profits. However, this is not always true: in Figure 3(b), the NPV does not consistently increase with income. This non-monotonicity when the interest rate is lower stems from the fact that on the one hand, the interest payments are lower and on the other hand, the present value of the default loss is higher with a higher $d$. Further, the lender employs a more aggressive lending strategy in the earlier periods, moving the expected default losses to these periods. The result is an NPV which first decreases and then starts to increase.
(a) $d=0.67$
(b) $d=0.83$
Figure 3: Lender NPV and its components as functions of the consumer type $k$ for uniformly-distributed income ($U[0,1]$) and $\rho=0.95$. In (a), $d=0.67$ and in (b), $d=0.83$.
Although the lender’s NPV is not necessarily monotone, the following Proposition shows that in addition to $\bar{x}$, which separates creditworthy from non-creditworthy consumers, there is a cutoff income value that separates profitable from unprofitable consumers. The resulting three (larger) segments are monotone in consumer income.
Proposition 2.6
Let income be uniformly distributed and let $k^{*}=\min\{k:\Pi_{k}\geq 0\}$. Then, unprofitable consumers default in periods $k=1,2,\ldots,k^{*}-1$; profitable consumers default in periods $k\geq k^{*}$; and creditworthy consumers never default. Further, there exists a cutoff value $\theta^{*}(\rho,d)=y_{k^{*}}$ such that
  • the consumer is unprofitable for $\theta\in[0,\theta^{*}(\rho,d))$,
  • the consumer is profitable (but non-creditworthy) for $\theta\in[\theta^{*}(\rho,d),\bar{x})$, and
  • the consumer is creditworthy for $\theta>\bar{x}$.
For all $k\geq k^{*}$, $\Pi_{k+1}\geq\Pi_{k}$.
The cutoff value $k^{*}$ is the minimum number of periods a consumer needs to repay her loans to be profitable. And, although the relationship between income and profitability is not necessarily monotone throughout, it is monotone among profitable (as well as creditworthy) consumers. Indeed, the relationship between NPV and consumer type is monotone for large enough $k$ in both Figures 3(a) and (b). The monotonicity of NPV for profitable customers reflects the fact that these consumers only become profitable in later periods. As outlined in section 2.2.1, in these later stages, the lender makes only small adjustments to the lending amounts, which results in a stable default loss. Once consumers keep borrowing for enough periods, their cumulative interest payments offset the final default loss.
2.3 Hybrid Information Architecture
One of the reasons online lending has attracted the attention of IS researchers is that it is enabled by modern IT that takes advantage of dynamic, real-time decision making (Hinz et al. (2011), Bharadwaj et al. (2013)). Online lending can change both the lending process and the information available to the lender and as such, provides an example of the effects of implementing modern IT architectures.
This is in contrast to the traditional credit evaluations performed by bank and credit card lenders before they underwrite a prospective borrower. Lenders traditionally evaluated credit applications upfront by examining the information available to them about the applicant from a variety of sources. They assessed applicants’ ability and willingness to repay based on their credit history and performance, employment history, past and outstanding obligations, cost of living, past transactions, and pre-existing loans vis-a-vis measures of income (FDIC (2007)). This information can be aggregated into a credit score that gives the lender a signal about the probability that the borrower would repay. The credit evaluation which led to the credit decision was conducted using a combination of human judgment and IT.
Traditional and direct lending are not mutually exclusive, however. In fact, a number of online DL, including international online DL Lenddo EFL, U.S. online DLs Upstart and Zest Finance (renamed Zest AI) and Mexican online DL MiMoni have become technology and process providers for traditional bank and credit union lenders, enabling them to use dynamic online direct lending as an add-on to their traditional underwriting processes. These DLs have merged their dynamic lending capabilities, AI prowess and understanding of social media and alternative information sources with the traditional lenders’ access to capital and customers to create hybrid lending architectures. Upstart, for example, uses AI models powered by more than 1,600 variables and more than 58 million past repayment events from its online lending experience to enable traditional lenders to lend online (Upstart Holdings, Inc. (2023)).
From an information acquisition perspective, the merger of traditional and online direct lending means that the lender has prior information about prospective borrowers as opposed to just having to rely on their repayment experience. We model a hybrid lender as one that combines online DL with a signal: in addition to knowing the income distribution $F$, the lender receives before period 00 a (one-time) signal $x_{0}$ which indicates that the consumer’s income is greater than or equal to $x_{0}$. After observing $x_{0}$, the lender decides on an initial loan amount $\mu_{0}$ (a function of $x_{0}$) which corresponds to repayment amount $y_{0}=(1+r)\mu_{0}=\mu_{0}/d$. If $y_{0}>\theta$, the consumer defaults; otherwise the lender receives the repayment amount $y_{0}$, updates the consumer income distribution, and the process continues.
[…]
Figure 4: Model Timeline, hybrid architecture.
Theorem 2.7 below shows that the optimal policy is in effect a hybrid policy which depends on the relationship between the signal $x_{0}$ and the lending threshold $\bar{x}$ of the dynamic LE policy. When $x_{0}\geq\bar{x}$, the lender will lend a constant amount $x_{0}$ in each period: if the prior information about the customer is sufficiently favorable, the lender opts for a static lending policy without any further information acquisition. However, when $x_{0}<\bar{x}$, the lender follows the structure of the LE strategy we found in the (pure) dynamic case (Theorem 2.4). In is case, the lender updates the population distribution $F$ to reflect the signal $x_{0}$.
Theorem 2.7
Optimal hybrid policy: there exists a unique optimal policy $\BFysuperscript\BFy\BFy^{*}$ which depends on $x_{0}$ and $\bar{x}$ as follows:
  • (a) If $x_{0}<\bar{x}$, the lender keeps increasing the repayment amounts $y_{t}$ in each period $t$ as long as the consumer repays her loans, and stops lending if the consumer defaults. The infinite sequence of (potential) repayment amounts $\{y_{t},t=0,1,2,\ldots\}$ is strictly increasing to the limit ${\bar{x}}$.
  • (b) If $x_{0}\geq\bar{x}$, the optimal policy is to set the same repayment amount $y_{t}=x_{0}$ in each period.
The hybrid policy is obviously more profitable than the dynamic policy as it takes advantage of additional information, but by how much, and what does its contribution depend on? First, it is straightforward that the benefit of the hybrid policy over the dynamic policy increases in $\rho$. Figure 5 compares the NPV of the hybrid and dynamic architectures for different values of the discount rate $d$. The Relative Advantage is defined as $\frac{E_{x_{0}}(\Pi_{h})-\Pi_{d}}{\Pi_{d}}$, where $E_{x_{0}}(\Pi_{h})$ represents the expected NPV under the hybrid architecture, taking the expectation over the signal distribution, with signal $x_{0}$ following the income distribution, and $\Pi_{d}$ denotes the expected NPV under the dynamic architecture.
Figure 5: NPVs of hybrid and dynamic architectures as functions of the discount rate $d$ for $\theta\sim U(0,1)$ and $\rho=0.95$.
As $d$ increases, both NPVs decrease because of a scale effect: As the interest rate decreases, lending is less profitable under both architectures, which also reduces the difference between the two NPVs. While for small $d$ the effect is almost linear in $d$, as $d$ approaches $\rho$, the lending amounts decline and the value of the signal relative to the value of dynamic lending information increases sharply. This happens because as $d$ increases towards $\rho$, the interest rate becomes low, which reduces the value of dynamic lending information relative to the value of the signal.
3 Endogenous Interest Rate
3.1 Optimal Information Acquisition with an Endogenous Interest Rate
3.1.1 Model and Optimal Policy
In many cases, the interest rate is set by the lender and becomes an endogenous decision variable. Within our model, this means that at the beginning of each period $t$, the lender sets both the repayment amount $y_{t}$ and the discount factor $d_{t}$ (the corresponding loan amount is $\mu_{t}=d_{t}y_{t}$). As before, the lender’s discount factor is $\rho\in(0,1)$. The consumer obviously prefers higher discount factors (lower interest rates) and her probability of accepting a loan from the lender increases with $d_{t}$. We model this using a demand function $s:[0,\rho]\to[0,1]$ that specifies for each discount factor $d\in(0,\rho)$ set by the lender, the probability that the consumer will accept the proposed loan terms; a borrower who rejects them leaves the system. We assume that the demand function is twice continuously differentiable, strictly increasing and strictly concave with $s(0)=0$, $s^{\prime}(\cdot)>0$ and $s^{\prime\prime}(\cdot)<0$.
The timeline is similar to our earlier model: the consumer’s income $\theta$ is drawn from a distribution $F$ with continuous positive support $[0,u)$. At $t=0$, the lender announces $(y_{0},d_{0})$, where $y_{0}$ is the initial repayment amount and $d_{0}$ is the initial discount rate. The consumer accepts the loan and borrows the offered amount with probability $s(d_{0})$. With probability $(1-s(d_{0}))$, the consumer rejects the loan and leaves the system. If $\theta<y_{0}$, the consumer defaults and leaves the system at $t=1$; if $\theta\geq y_{0}$, the consumer repays $y_{0}$, the lender announces $(d_{1},y_{1})$, and the process continues.
[…]
where we define $\Pi^{-1}_{s=0}\operatorname{\mathbb{P}}(\theta\geq y_{s}|\mathcal{H}_{s})s(d_{s})=1$.
Let the state of the system be $x_{t}=\max_{i=0\dots,t-1}\{y_{i}\}$, the current lower bound on $\theta$. Then, $x_{t+1}=\max(x_{t},y_{t})$ and the Bellman equation is
${J(x_{t})\ =\max_{y_{t},d_{t}}\ \{-s(d_{t})d_{t}y_{t}+\rho s(d_{t})\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})(J(x_{t+1})+y_{t})\}.}$ (11)
The counterpart of Lemma 2.2 in this case is:
Lemma 3.1
If an optimal policy $\{y_{t},d_{t}\}$ exists, then $y_{t}$ is weakly increasing, i.e., $y_{t}\geq y_{t-1}$ for all $t>0$.
The intuition is similar to that of Lemma 2.2: when the lender knows that $\theta\geq y_{t-1}$, setting $y_{t}<y_{t-1}$ is sub-optimal. Thus, the state evolves as $x_{t}=y_{t-1}$ and the Bellman equation becomes
${J(x_{t})\ =\ \max_{y_{t},d_{t}}\ \ \{x_{t}-s(d_{t})d_{t}y_{t}+\rho s(d_{t})\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})J(y_{t})\}.}$ (12)
Lemma 3.2
There exists a unique finite solution $0<\bar{x}<u$ to the equation:
${G(\bar{x})=-\frac{(1-s(d^{*})\rho)}{(1-s(d^{*})d^{*})}\frac{d^{*}}{\rho}+1,}$ (13)
where $d^{*}\in(0,\rho)$ is the unique solution of
${\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}=\frac{s(d^{*})+s^{\prime}(d^{*})d^{*}}{\rho s^{\prime}(d^{*})}.}$ (14)
Moreover, when $s(d)=d^{\alpha}$ with $\alpha\in(0,1)$, $d^{*}$ is increasing in $\alpha$.
The structure of the optimal policy is largely driven by the monotonicity properties of the demand elasticity (with respect to the discount rate), given by
${\xi(d)=\dfrac{ds^{\prime}(d)}{s(d)}.}$ (15)
In particular, when the demand elasticity is monotone decreasing, the structure of the optimal lending policy is similar to the one we found in Theorem 2.4:
Theorem 3.3
Optimal policy: When $\xi(d)$ is monotone decreasing, there exists a unique optimal policy whereby the lender keeps increasing the repayment amounts $y_{t}$ in each period $t$ as long as the consumer repays her loans. The infinite sequence of repayment amounts $\{y_{t},t=0,1,2,\ldots\}$ is strictly increasing to the limit ${\bar{x}}$. The optimal discount rates are not necessarily monotone and converge to $d^{*}$.
Intuitively, when the demand elasticity is decreasing, the consumer is less likely to balk at higher discount rates. The discount rate generally increases (interest rate decreases) as the lender accumulates more positive information about the consumer’s creditworthiness. This creates an advantage for experimenting later, when the discount rate is higher. Thus, the lender’s optimal policy is based on an infinite number of small experiments, as aggressive early experiments will reduce demand by all consumers, leading to the loss of too many creditworthy ones. This is an LE policy similar to the one we obtained with an exogenous interest rate. However, when the demand function has constant elasticity, $s(s)=d^{\alpha}$ with $\alpha\in(0,1)$, the optimal policy structure changes dramatically.
Theorem 3.4
Optimal policy: When the demand elasticity is constant, there exists a unique optimal policy characterized by a single grand experiment (GE): the lender first sets the repayment amount at $y_{0}=\bar{x}$ and the discount rate at $d_{0}=(1-F(\bar{x}))d^{*}$. If the consumer repays $y_{0}$, the lender fixes the repayment amount at $y_{t}=\bar{x}$ and the discount rate at $d_{t}=d^{*}$ for all subsequent periods.
Theorem 3.4 is remarkable, showing that when the demand elasticity is constant, the lender designs one GE to test upfront whether the consumer can repay the optimal long-term repayment amount $\bar{x}$. If the answer is positive, it’s unnecessary to take any further information acquisition risk and the lender charges in each period the same interest rate and lends in each period the same amount which together lead to repayment amount $\bar{x}$. From period 1 on, the lender takes no risk: it knows the consumer can repay $\bar{x}$. Further, these interest rate and loan amount are at their long-term optimal levels. The lender takes advantage of its ability to control both the discount rate and the loan amount when it designs the grand experiment by setting a lower loan amount and a higher interest rate that together lead to the optimal repayment amount $\bar{x}$. In this way, one test establishes everything the lender needs to know without taking excessive risk.
The GE strategy does not hinge on the constant elasticity assumption, although the case where demand has increasing elasticity is more complex.
Theorem 3.5
Optimal policy: When both the demand elasticity and the hazard rate function $\frac{f(x)}{1-F(x)}$ are increasing, there exists a unique optimal policy. Under the optimal policy, the lender sets the initial repayment amount $y_{0}\geq\bar{x}$ and the initial discount rate $d_{0}$ to jointly solve
${G(y_{0})+\dfrac{d_{0}}{\rho(1-F(y_{0}))}\dfrac{(1-s(d^{*})\rho)}{(1-s(d^{*})d^{*})}=1,\quad\dfrac{(1-F(y_{0}))(1-s(d^{*})d^{*})}{(1-s(d^{*})\rho)}=\dfrac{s(d_{0})+s^{\prime}(d_{0})d_{0}}{\rho s^{\prime}(d_{0})}.}$ (16)
If the consumer repays $y_{0}$, the lender fixes the repayment amount at $y_{t}=y_{0}$ and discount rate at $d_{t}=d^{*}$ for all subsequent periods.
Under the conditions of Theorem 3.5, the lender still employs a single GE. However, the experiment does not involve repayment amount $\bar{x}$. Instead, the lender tests whether the consumer can repay a higher amount $y_{0}$. The reason is that with increasing demand elasticity, the consumer is least sensitive to a higher interest rate when it’s high. Taking advantage of that, the lender designs a more aggressive GE with a higher repayment amount than $\bar{x}$. Similar to the constant elasticity case, the first loan has a lower discount rate than $d^{*}$ and if the consumer repays $y_{0}$, the lender will not take any risk in subsequent periods by lending $(d^{*}y_{0})$ in each period and setting the discount rate at its optimal level, $d^{*}$.
3.2 Implications
3.2.1 LE vs. GE
Theorems 2.4 and 3.4 show that under our dynamic model, either LE or GE approach may be optimal depending on the properties of the demand function and the scope of the lender’s decisions (endogenous vs. exogenous interest rate). Lenders should adopt a LE policy when the demand is entirely inelastic (exogenous rate) or when the demand elasticity is decreasing in the discount rate. They should follow GE when the demand elasticity is constant or increasing in the discount rate. The distinction between the two cases lies in the lender’s consideration of the risk of losing customers when it charges a high interest rate (low discount rate), which hinges on the shape of the demand function $s(d)$.
The intuitive reason this happens is that the lender has to balance its default risk from low-income borrowers against the risk of losing profitable high-income borrowers. Early on, the lender will generally charge a high interest rate (low discount rate) due to the high risk of default. Once the customer pays back, the lender tends to lower its interest rate to retain that customer. If the demand elasticity is increasing, the higher risk of losing more profitable customers pushes experimentation to the earliest possible period, resulting in one early GE. If the demand elasticity is decreasing, later experiments entail a lower risk of losing customers, and coupled with the lower default losses associated with smaller experiments, this implies that the LE policy is optimal.
The implications of the LE strategy studied in Section 2.2.1 carry over qualitatively to the decreasing demand elasticity case with endogenous rates. We have also seen (Theorems 3.4 and 3.5) that both the constant and increasing elasticity cases lead to qualitatively similar GE results. Hence, we study next the implications of the constant elasticity case, which is more transparent, as a canonical GE case.
3.2.2 Consumer Segmentation under GE
When the lender employs a single GE, the consumer can default only when $\theta$ is less than the first repayment amount, $\bar{x}$ (in the increasing elasticity case, $\bar{x}$ is replaced by $y_{0}$). This gives rise to two consumer segments:
  • Unprofitable consumers with $\theta<\bar{x}$, who default in the first period. The default results in a loss to the lender.
  • Creditworthy consumers with $\theta\geq\bar{x}$, who repay in the first period. These consumers are profitable.
Unlike the LE case, un-creditworthy consumers cannot be profitable, and consumers’ NPV is always an increasing function of their income $\theta$.
3.3 Hybrid Information Architecture
As we have done in the exogenous interest rate case, we next construct a hybrid model which starts with an information signal before $t=0$ and continues with a dynamic lending policy. We provide here the optimal policy for the hybrid architecture in the constant elasticity case (the other cases are analyzed in detail in Appendices A.9 for decreasing elasticity and A.11 for increasing elasticity).
Theorem 3.6
Optimal hybrid policy: Under the hybrid architecture with constant demand elasticity, there exists a unique solution to the Bellman equation given by
${J(x_{t})=\begin{cases}x_{t}-(d_{x_{t}})^{\alpha+1}\bar{x}+\rho(d_{x_{t}})^{\alpha}\dfrac{1-F(\bar{x})}{1-F(x_{t})}\dfrac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}\bar{x},\quad x_{t}<\bar{x},\\ \dfrac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}x_{t},\quad x_{t}\geq\bar{x},\end{cases}}$ (17)
where
${d_{x_{t}}=\frac{1-F(\bar{x})}{1-F(x_{t})}d^{*}and\quad\bar{x}=G^{-1}\left(\frac{1}{\alpha+1}\right).}$ (18)
[…]
  • (a) if $x_{0}<\bar{x}$, the lender sets the repayment amount at $y_{0}=\bar{x}$ with the corresponding discount rate $d_{0}=\frac{1-F(\bar{x})}{1-F(x_{0})}d^{*}$. If the consumer repays $y_{0}$, the lender fixes the repayment amount $y_{t}=\bar{x}$ and discount rate $d_{t}=d^{*}$ for all future periods.
  • (b) if $x_{0}\geq\bar{x}$, the optimal policy is to set $y_{t}=x_{0}$ in each period with a fixed discount rate $d^{*}$.
Theorem 3.6 shows, again, that the hybrid architecture leads to a hybrid policy which combines a static lending amount for richer consumers with a GE policy for poorer customers. When the signal indicates that the consumer has income below the lending threshold $\bar{x}$, the lender employs a GE strategy with a single experiment designed to test whether the customer can repay $\bar{x}$. When the signal indicates that the customer’s income is above $\bar{x}$, the lender foregoes the opportunity to experiment and lends the lower bound of the consumer’s income in all periods, similar to what we have seen in the LE case.
As before, these results extend in a natural way to the case of increasing elasticity and hazard rate, with the policy structure being a hybrid of GE and a constant lending amount. There is one nuanced difference: when the elasticity is increasing in the discount rate, the lender benefits from a more aggressive grand experiment. As a result, the repayment amount becomes $y_{0}\geq\bar{x}$. We provide the complete analysis of this case in appendix A.11.
With the optimal policy in hand, we can now compare the performance of the hybrid and dynamic architectures under GE. Let $\Pi_{h}(x_{0})$ be the lender’s expected NPV under the hybrid policy when its signal is $x_{0}$. Since the Bellman function $J(x_{t})$ defined in (17) is the sum of the current repayment amount $x_{t}$ and the future expected NPV under $x_{t}$, the expected NPV with signal $x_{0}$ is given by $\Pi_{h}(x_{0})=J(x_{0})-x_{0}$. As before, the benefit is increasing in the lender’s discount rate $\rho$.
Figure 6: NPVs of hybrid and dynamic architectures as functions of $d^{*}$ for the uniform income distribution with $\rho=0.95$.
Figure 6 compares the NPVs of the hybrid and dynamic architectures as functions of $d^{*}$. The Relative Advantage (of the hybrid architecture) is defined as $\frac{E_{x_{0}}(\Pi_{h})-\Pi_{d}}{\Pi_{d}}$, where simiarly, the expectation is taking over the signal distribution, with signal $x_{0}$ following the same distribution as consumer income. Using equation (14), for each value of the demand elasticity $\alpha$ we calculated the corresponding long-run discount factor $d^{*}$ (an increasing function of $\alpha$, as in Lemma 3.2) and plotted the corresponding NPVs. In this way, we can compare the GE results to those previously obtained under LE (Figure 5). First, the endogenous $d^{*}$ never reaches $\rho$ because it is suboptimal to offer very low interest rates. Second, the NPVs decline less sharply as $d^{*}$ increases and the advantage of the hybrid architecture is larger for the feasible values of $d^{*}$. As $\alpha$ and $d^{*}$ increase, the absolute advantage of the hybrid architecture decreases (a scale effect similar to the LE case) but its relative advantage increases.
3.4 Inclusive Finance
The simple policy structure of GE (in particular, the constant elasticity case) allows us to examine the effects of the hybrid architecture on financial inclusion. In a review of inclusive finance, Corrado and Corrado (2017) refer to inclusive finance as the provision of loans and other basic financial services, often through microfinance institutions, to those often left out, so that “access to financial services, primarily to credit, can enable agents to make longer-term consumption and investment plans, participate in productive economic activities and cope with adverse shocks” (Corrado and Corrado (2017)). Multiple world agencies from the United Nations to the World Bank and the International Monetary Fund sponsor initiatives that support inclusive finance to make poorer people better off (see, for example, Kara et al. (2021)). IT plays a key role in achieving these maxims (Balasubramanian (2022), McKinsey Global Institute (2016), Lapukeni (2015), Gupta and Kanungo (2022)), an effect confirmed by multiple empirical papers. Balyuk (2023) shows that Prosper contributed to financial inclusion with a spillover effect on US banks, and Gopal et al. (2023, 2024) show using LC data that their portfolio construction methodology improves financial inclusion while making loan portfolios more attractive to institutional investors. Malladi et al. (2021) and Kshetri (2021) show how the Digital India initiative and more broadly, the use of ICT drove financial inclusion in India, Girma and Huseynov (2023) establishe the relationship in Africa, Chu et al. (2016) show how taking major lender CreditEase online contributed to inclusive finance in China, a World Bank report (Pazarbasioglu et al. (2020)) surveys similar initiatives around the world, and Afjal (2023) provides a bibliometric survey.
Our model shows that the hybrid lending architecture has important advantages over the dynamic architecture in terms of lender profitability. Further, it makes richer consumers better off by lending them $x_{0}>\bar{x}$ whereas the loans made to poorer consumers are bounded by $\bar{x}$. Does this happen at the expense of poorer consumers? Is the hybrid architecture less inclusive than the dynamic architecture?
We now show that this is not the case. In fact, the hybrid architecture makes both poorer and richer consumers better off. To show this pareto-improvement, we define a strong form of inclusion based on stochastic dominance (Levy (1992)). We consider two forms of inclusion: by coverage and by loan amounts. We say that architecture A is more inclusive than architecture B by coverage if for each consumer and in each time period, the probability that the consumer gets a loan under architecture A is greater than or equal to the probability that the consumer gets a loan under architecture B. We say that architecture A is more inclusive than architecture B by loan amounts if for each consumer and in each time period, the loan amounts offered to the consumer under architecture A stochastically dominate the loan amounts offered under architecture B. These are strong requirements which are hard to satisfy. However, we prove that in the constant elasticity case studied here, the hybrid architecture is more inclusive than the dynamic architecture by both coverage and loan amounts.
Proposition 3.7
In the constant elasticity case, the hybrid architecture is more inclusive than the dynamic architecture by both coverage and loan amounts.
In terms of loan amounts, the signal $x_{0}$ information enables the hybrid architecture to lend more. In terms of coverage, both architectures offer loans to all consumers in the initial period. Subsequently, the hybrid architecture has a lower probability of default and a higher or equal probability of accepting the loan than the dynamic architecture. As a result, the retention probabilities of the hybrid architecture are higher, implying that it is more inclusive by coverage. Figure 7 compares the loan amounts and the initial period retention probability as functions of $x_{0}$ under the two architectures.
(a) Loan amount
(b) Initial period retention probability
Figure 7: Loan amount and initial period retention probability as functions of $x_{0}$ under dynamic and hybrid architectures with $\alpha=0.5$, $\rho=0.95$, and uniform income distribution.
As shown in Figure 7, for each $x_{0}$, the hybrid lender lends more than the dynamic lender in the initial period. In all subsequent periods, the hybrid lender lends the same amount as the dynamic lender when $x_{0}\leq\bar{x}$ and more when $x_{0}>\bar{x}$. Further, under the hybrid architecture, the retention probability following the initial period is higher than that under the dynamic architecture, and the same applies to all subsequent periods. It follows that the hybrid architecture is more inclusive: it offers larger (or the same) loan amounts to all consumers, and it offers loan to more of them. In this case, what is good for the poor is also good for the rich, as well as for the lender.
4 Income variability
We study the effects of income variability on lender profitability for symmetric Beta-distributed income with density $f(a,b,x)=\frac{x^{a-1}(1-x)^{b-1}}{B(a,b)}$ where $B(a,b)=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}$ and $a=b$, so the mean income is $\frac{1}{2}$ ($\Gamma(\cdot)$ is the Gamma function). Decreasing $a$ increases the variance of income, which is given by $\frac{1}{4(2a+1)}$. One intuitively expects lenders to prefer lower income variability over higher variability. Figure 8, however, shows this is not the case: lender profitability is not monotone in the income variance, and a higher variance may actually benefit the lender. In fact, the following Proposition shows that for the symmetric Beta distribution, the lender’s NPV is always U-shaped.
[…]
Proposition 4.1
For the symmetric Beta distribution with a fixed mean, the lending profit first decreases and then increases as the income variance increases.
The common decreasing relationship between NPV and variance reflects lender uncertainty. With no income variance, the lender has full information and can set the repayment amount equal to the consumer’s income without any losses. As the income variance increases, the lender may lose money and to protect itself, it lends less. Its profits decline as a result of both its expected losses and its lower interest income.
The increasing part of the U-shape reflects the option value of experimentation. As the variance increases, the option value of the grand experiment increases as it increases the likelihood of identifying high-income borrowers. This also implies that the lender would increase the lending threshold in order to better target high-income borrowers. This effect has to be balanced against the usual negative effect of uncertainty. The option value is particularly high for high values of the income variance, which leads to the U shape.
We illustrate these general considerations with a simple example. Consider a two-point income distribution with a fixed mean, $u$, where $\theta$ is $u+\delta$ or $u-\delta$ with probability $\frac{1}{2}$ each. In the special case where $u=\delta=\frac{1}{2}$, the income distribution is Beta with $a=b=0$; when $u=\frac{1}{2}$ and $\delta=0$, the income distribution is deterministic—a limiting form of the Beta distribution with $a$ and $b$ going to infinity.
It is easy to see that one of following two policies must be optimal:
  1. (A)
    At $t=0$, lend $(u-\delta)+\epsilon$ at discount rate $d_{0}$ with $\epsilon$ being a small positive number, and proceed to offer $(u+\delta)$ in each period if the first-period loan is repaid. The corresponding expected NPV as $\epsilon\to 0$ is $\Pi^{A}(\delta)=-(d_{0})^{a+1}(u-\delta)+\frac{1}{2}\rho(d_{0})^{\alpha}\left((u-\delta)+\left(\frac{1-(d^{*})^{\alpha+1}}{1-\rho(d^{*})^{\alpha}}-1\right)(u+\delta)\right)$.
  2. (B)
    Lend $(u-\delta)$ with discount rate $d^{*}$ in each period unless the borrower defaults. The corresponding expected NPV is $\Pi^{B}(\delta)=\left(\frac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}-1\right)(u-\delta)$.
For each $\delta$, the lender’s expected profit is $\max\left\{\Pi^{A}(\delta),\Pi^{B}(\delta)\right\}$, the maximum of two linear functions of $\delta$, where increasing $\delta$ increases the variance of the income distribution. Under Policy (A), the first experiment gives the lender the option to identify the high-income consumer. As is common in option valuation, the option value is an increasing function of the variance (parameterized here by $\delta$) – if the consumer has the low income $u-\delta$, she will default and won’t get another loan anyway. Policy (B) lends to the entire market, including low-income consumers, and its profit decreases in $\delta$ (i.e., profits decrease with variance, as usual), as it does not allow the lender to identify the high-income consumer and as a result has no option value. Rather, under policy (B) the lender “plays it safe.” For large enough $\delta$ (and variance), policy (A), which targets the high income consumer, yields a higher expected profit. For small $\delta$, the income differences are small and the payoff of experimentation is low. As a result, for small $\delta$ the lender is better off with policy (B) which lends less to the entire market. The linear profit functions $\Pi^{A}(\delta)$ and $\Pi^{B}(\delta)$ intersect at some $\delta^{*}$, so the expected profit function $\max\left\{\Pi^{A}(\delta),\Pi^{B}(\delta)\right\}$ is U-shaped (in this case, V-shaped) with a minimum at the intersection point $\delta^{*}$.
Similar arguments hold in the LE case, where again, numerically, the expected NPV is a U-shaped function of the income variance.
5 Concluding Remarks
Developments in ICT gave rise to online lending, which has transformed lending and attracted significant interest in the IS literature. This paper examines the optimal information acquisition strategy for a direct online lender that provides unsecured loans to consumers with unknown incomes. We present optimal policies for models with exogenous and endogenous interest rates. In the exogenous rate scenario, the lender gradually increases the repayment amount, converging to a threshold; this solution corresponds to a Lean Experimentation paradigm. Under endogenous rates, the shape of the demand elasticity function determines the structure of the optimal policy. When the demand elasticity decreases, LE is still optimal. When the demand elasticity is constant or increasing, the lender engages in a single Grand Experiment without investing in information acquisition in any subsequent period. We derive the optimal policies explicitly and (among other results) provide sharp predictions on the effects of income variability, the structure of consumer segmentation, and the effects of hybrid information architectures on lender profitability and financial inclusion.
Our model is stylized and we believe an important next step is to subject our results to empirical validation. The structure of lending policies and their impact on profitability may be tested directly using data from direct lenders. Our predicted effects of the transition to hybrid architectures may similarly be tested directly using lender data. Such data are not trivial to acquire, but a number of IS researchers have managed to use comparable datasets in other contexts. A less challenging feat is to test our predictions using macro data. Online lending has been adopted in multiple countries, and cross-country data may be used to test some of our predictions, including the drivers of LE vs. GE strategies, the growth of hybrid architectures, the effects of income variability and the effect of hybrid lending on financial inclusion. For example, our models predict that highly regulated interest rate environments lead to the adoption of LE strategies whereas flexible rate environments lead (under the conditions studied in Section 3.1) to GE strategies. Thus, regulatory differences among states and countries may lead to sharply different lending strategies. Regarding income variability, our models predict the counter-intuitive result that the relationship between income variance and profitability or ultimate lending amounts (represented in our model by the limit $\bar{x}$) is not monotone. Such relationship may be tested using income variances as proxies for income uncertainty. Regarding financial inclusion, in Section 3.4 we cited multiple studies that tested the effects of different aspects of Fintech on financial inclusion using a number of dependent variables. These methodologies may be adapted to test the effects of hybrid lending on financial inclusion.
Although this requires further study, we expect some of the insights we derived for the online lending market to extend to other settings. At a minimum, we have shown that the commonly-assumed lean experimentation strategy is not optimal for an important sector of the economy. While LE is optimal in repeated natural experiments, where the experiment does not change the underlying probabilities, experimenting with customers changes their behavior and calls for more nuanced models (in our model, the properties of the demand elasticity drove the choice between GE and LE). Further, we found that multi-dimensional experiments (in our model, involving both interest rates and loan amounts) allow decision-makers to increase the power of a single experiment to the point where they can derive sufficient statistics for all subsequent periods. Our intuition is that other settings that share similar properties will give rise to information acquisition strategies that are limited to one or a small number of experiments. (This will of course happen when there is a high fixed cost per experiment, but we are interested in settings where the result arises more naturally).
One interesting case which we think deserves further study is information acquisition in the process of systems development. As discussed earlier, Continuous Deployment (CD) enables systems developers to engage in a large number of small information acquisition experiments through automated testing and delivery of systems directly to the production environment. An interesting question which may combine theoretical and empirical research is: under what conditions is CD optimal? One driving factor is industry clockspeed (Mendelson and Pillai (1998)). Controlling for industry clockspeed, however, which other firm and industry characteristics drive the adoption of CD? More broadly, what types of firms and industries would prefer the LE approach and which ones would prefer GE?
A Proofs
A.1 The Gamma, Weibull and Beta distributions satisfy Assumption 2.1
A.1.1 The Gamma Distribution satisfies Assumption 2.1
Consider the two-parameter Gamma distribution with shape parameter $a>0$, scale parameter $b>0$, and probability density function (pdf) $f(x)=\frac{b^{a}}{\Gamma(a)}x^{a-1}e^{-bx}$. Then,
[…]
It’s straightforward that $\lim_{x\to 0}G(x)=0$. For $\lim_{u\to\infty}G(u)\geq 1$, substituting $y=t-x\geq 0$, we have
[…]
A.1.2 The Weibull Distribution satisfies Assumption 2.1
Consider the two-parameter Weibull distribution with shape parameter $k>0$, scale parameter $\lambda>0$, and pdf $f(x)=\frac{k}{\lambda}\left(\frac{x}{\lambda}\right)^{k-1}e^{-(x/\lambda)^{k}}$. Then, $G(x)=xH(x)$ is given by
${G(x)=\frac{xf(x)}{1-F(x)}=\frac{\frac{k}{\lambda}\left(\frac{x}{\lambda}\right)^{k-1}e^{-(x/\lambda)^{k}}}{e^{-(x/\lambda)^{k}}}=k\left(\frac{x}{\lambda}\right)^{k},}$ (20)
which implies that $\lim_{x\to 0}G(x)=0$, $\lim_{x\to\infty}G(x)=\infty$, and for $k,x>0$, $G(x)$ is strictly increasing. $\square$
A.1.3 The Beta Distribution satisfies Assumption 2.1
Consider the Beta distribution with positive parameters $a$ and $b$ and pdf ${f(x)=\frac{1}{B(a,b)}x^{a-1}(1-x)^{b-1},}$ where $B(a,b)=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}$ and $\Gamma(\cdot)$ is the Gamma function. Proving that this distribution satisfies Assumption 2.1 is more complex. We first show that for all $a,b>0$, $G(0)=0$ and $\lim_{x\to 1}G(x)\geq 1$. We then show that $G(x)$ is strictly increasing.
It’s straightforward that $G(0)=0$ for all $a,b>0$. To show that $\lim_{x\to 1}G(x)\geq 1$, we consider the two cases (i) $b\leq 1$ and (ii) $b>1$. (i) When $b\leq 1$, $f(1)>0$ and as ${x\to 1}$, $\frac{xf(x)}{1-F(x)}>1$ as it goes to infinity. (ii) When $b>1$,
[…]
To show that $G(x)$ is increasing, we consider three cases:
  • Case 1: $a\geq 1$. It is well known that for $a\geq 1$, the Beta distribution has an increasing failure rate, which implies that $G(x)$ is increasing.
  • Case 2: $a<1$ and $b\leq 1$. The derivative of $G(x)$ is ${G^{\prime}(x)=\frac{(f(x)+xf^{\prime}(x))(1-F(x))+xf(x)^{2}}{(1-F(x))^{2}}.}$ The sign of $G^{\prime}(x)$ is the same as ${xf^{\prime}(x)+f(x)=\frac{a(1-x)+(1-b)x}{B(a,b)x^{1-a}(1-x)^{2-b}}}$. Since $b\leq 1$, $xf^{\prime}(x)+f(x)>0$ for all $x\in(0,1)$, hence $G^{\prime}(x)>0$.
  • Case 3: $a<1$ and $b>1$. Define the inverse hazard function $g(x)=\frac{1}{H(x)}=\frac{1-F(x)}{f(x)}$, which is bathtub-shaped in this case (Glaser [1980]). Thus, there is a unique $x_{0}\in(0,1)$ such that $g^{\prime}(x)>0$ for $x<x_{0}$, $g^{\prime}(x_{0})=0$ and $g^{\prime}(x)<0$ for $x>x_{0}$. Now, $G(x)=xH(x)$ is increasing if and only if its inverse $\frac{1}{G(x)}=\frac{1-F(x)}{f(x)x}=\frac{g(x)}{x}$ is decreasing, i.e., $\left[\frac{1}{G(x)}\right]^{\prime}=\frac{g^{\prime}(x)x-g(x)}{x^{2}}<0$, which holds if the numerator $\xi(x)=g^{\prime}(x)x-g(x)$ is negative. We show this is the case by considering the two ranges: $x\in[x_{0},1]$ and $x\in[0,x_{0})$.
    For $x\in[x_{0},1]$, $g^{\prime}(x)\leq 0$ and $g(x)>0$, hence $\xi(x)=xg^{\prime}(x)-g(x)<0$.
    For $x\in[0,x_{0})$, define $\eta(x)=-\frac{f^{\prime}(x)}{f(x)}=\frac{x(a+b-2)+(1-a)}{x-x^{2}},$ so
    ${g^{\prime}(x)=g(x)\eta(x)-1.}$ (21)
    Now, ${\eta^{\prime}(x)=\frac{a-1}{x^{2}}+\frac{b-1}{(1-x)^{2}}\ \textup{and}\ \eta^{\prime\prime}(x)=2\left(\frac{1-a}{x^{3}}+\frac{b-1}{(1-x)^{3}}\right).}$ With $\eta^{\prime\prime}(x)>0$ for all $x\in(0,1)$, $t_{0}=\frac{1}{1+\sqrt{(b-1)/(1-a)}}\in(0,1)$ is the unique global minimum of $\eta(x)$ satisfying $\eta^{\prime}(x)=0$. Further, since $\eta(t_{0})>0$, $\eta(x)$ is positive for all $x\in(0,1)$.
    With these preliminaries, we now prove that $\xi(x)=xg^{\prime}(x)-g(x)<0$ for $x\in(0,x_{0})$. Rewrite $\xi(x)$ as
    ${\xi(x)=xg^{\prime}(x)-g(x)=xg^{\prime}(x)-\frac{g^{\prime}(x)+1}{\eta(x)}=\frac{g^{\prime}(x)\left(\eta(x)x-1\right)-1}{\eta(x)},}$ (22)
    where $g(x)=\frac{g^{\prime}(x)+1}{\eta(x)}$ by (21). As $\eta(x)>0$, showing that $\xi(x)<0$ is equivalent to showing that the numerator $\gamma(x)$ of equation (22), given by
    ${\gamma(x)=g^{\prime}(x)(\eta(x)x-1)-1=g^{\prime}(x)\left(\frac{(a+b-1)x-a}{1-x}\right)-1,}$ (23)
    is negative. To that end, we consider two ranges: (i) $x\in(0,\min\{\frac{a}{a+b-1},x_{0}\}]$, and (ii) $x\in(\min\{\frac{a}{a+b-1},x_{0}\},x_{0})$ (which may be empty). For (i), $\eta(x)x-1\leq 0$ and $g^{\prime}(x)>0$, so $\gamma(x)$ is negative. For a nonempty (ii), substitute $g^{\prime}(x)$ from (21):
    ${\gamma(x)=\eta(x)(g(x)(\eta(x)x-1)-x).}$ (24)
    For $x>\frac{a}{a+b-1}$, $\eta(x)x-1=\frac{(a+b-1)x-a}{1-x}>0$. As $x_{0}$ is the maximizer of $g(x)$, ${g(x)(\eta(x)x-1)<g(x_{0})(\eta(x)x-1).}$ At $x_{0}$, $g^{\prime}(x_{0})=g(x_{0})\eta(x_{0})-1=0$, so $g(x_{0})=\frac{1}{\eta(x_{0})}$. From equation (24),
    ${\gamma(x)<\eta(x)\left(g(x_{0})(\eta(x)x-1)-x\right)=\eta(x)\left(\frac{x}{\eta(x_{0})}[\eta(x)-\frac{1}{x}-\eta(x_{0})]\right).}$
    Let $\zeta(x)=\eta(x)-\frac{1}{x}$. For $a<1$ and $b>1$, $\zeta(x)=\frac{(x(a+b-1)-a)}{x(1-x)}$ and $\zeta^{\prime}(x)=\frac{a(1-x)^{2}+(b-1)x^{2}}{(x^{2}(1-x)^{2})}>0,$ so $\zeta(x)$ is monotone increasing. Now, $\zeta(x)<\zeta(x_{0})$ for $x\in(\frac{a}{a+b-1},x_{0})$, hence $\xi(x)<\xi(x_{0})<\eta(x_{0})$ and
    ${\gamma(x)<\eta(x)\left(\frac{x}{\eta(x_{0})}[\xi(x)-\eta(x_{0})]\right)<\eta(x)\left(\frac{x}{\eta(x_{0})}[\xi(x_{0})-\eta(x_{0})]\right)<0,}$
which completes the proof. $\square$
A.2 Proof of Lemma 2.2
For any policy, let $t$ be the first time $y_{t}<y_{t-1}$, and let $x_{t}=\max_{-1\leq i<t}y_{i}=y_{t-1}$. We compare the expected NPV from period $t$ onwards for setting $y_{t}<y_{t-1}$ versus deviating to $\tilde{y}_{t}=y_{t-1}$. Note that under both policies, the state in period $t+1$ is $x_{t+1}=\max_{-1\leq i<t+1}\ y_{i}=x_{t}$, so the future continuation value $J(x_{t})$ is the same under both policies.
If the lender sets $y_{t}<x_{t}$, its expected NPV will be ${-dy_{t}+\rho\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})[y_{t}+J(x_{t})]=(\rho-d)y_{t}+\rho J(x_{t}),}$ since $\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})=1$. The NPV with $\tilde{y}_{t}=y_{t-1}=x_{t}$ is ${\left(\rho-d\right)x_{t}+\rho J(x_{t}).}$ The difference between the two NPV values is $\left(\rho-d\right)(x_{t}-y_{t})>0$. Since this inequality holds for all $t$, it’s never optimal to set $y_{t}<x_{t}$. $\square$
A.3 Proof of Lemma 2.3
Define $T(x)=G(x)-\delta$ where $\delta=\frac{\rho-d}{\rho(1-d)}\in(0,1)$. By assumption 2.1, $T(0)<0$, $\lim_{x\to u}T(x)>0$, and $T(\cdot)$ is strictly increasing in the interval $[0,u)$. It follows from the continuity of $T(\cdot)$ that there exists a unique root $\bar{x}$ of $T(x)=0$ (which is equivalent to (4)). Since $G(\cdot)$ is increasing, its inverse $G^{-1}(\cdot)$ exists and is increasing. Now, $\bar{x}=G^{-1}(\delta)$, and the rest of the proof follows from the fact that $\delta$ is increasing in $\rho$ and decreasing in $d$. $\square$
A.4 Proof of Theorem 2.7 (and Theorem 2.4)
Theorem 2.4 is a special case of Theorem 2.7, so we prove the latter. The proof is structured as follows. We first establish the existence and uniqueness of the solution to the Bellman equation (Lemma A.1). Next, Lemma A.2 helps us establish that the optimal repayment amount must follow one of three structures: (1) constant amount, (2) strictly increasing amounts, or (3) initially strictly increasing followed by a constant amount. We then build the proof for the two parts of the Theorem, (a) $x_{0}<\bar{x}$ and (b) $x_{0}\geq\bar{x}$.
Lemma A.1
There exists a unique solution to the Bellman equation (3).
Proof. First, we can bound the repayment amounts $y_{t}$ for all $t$ since in any period $t$, there is an upper bound on $y_{t}$ above which the expected payoff becomes negative (in fact, as $y_{t}\to\infty$, the expected payoff $\lim_{y_{t}\to\infty}x_{t}-dy_{t}\to-\infty$). Thus, for some $M>0$ we can optimize (3) within $y_{t}\in[0,M)$ for all $t$. With bounded controls and discount factor $\rho\in(0,1)$, there exists a unique solution to the Bellman equation (see, e.g. Bertsekas [2012]). $\square$
Building on the monotonicity result of Lemma 2.2, we next show that if the optimal control satisfies $y_{t}=y_{t-1}=x_{t}$, then the optimal policy sets $y_{k}=x_{k}$ for all subsequent periods.
Lemma A.2
[…]
Proof. If $y^{*}_{t}=x_{t}$, then $x_{t}$ maximizes the following Bellman equation:
${J(x_{t})\ =\ \max_{y_{t}\in[x_{t},u)}\ \ {x_{t}-dy_{t}+\rho\dfrac{1-F(y_{t})}{1-F(x_{t})}J(y_{t})}.}$ (25)
By Lemma 2.2, the law of motion evolves as $x_{t+1}=y_{t}=x_{t}$, so the Bellman equation for period $t+1$ is the same as for period $t$:
${J(x_{t+1})=J(x_{t})\ =\ \max_{y_{t+1}\in[x_{t},u)}\ \ {x_{t}-dy_{t+1}+\rho\dfrac{1-F(y_{t+1})}{1-F(x_{t})}J(y_{t+1})}.}$ (26)
As $x_{t}$ maximizes (25), it also maximizes (26), so $y^{*}_{t+1}=y^{*}_{t}=x_{t}$. The Lemma now follows by forward induction. $\square$
It follows that there can only be three possible structures for the optimal policy: (1) $\{y_{t},\ t\geq 0\}$ are constant; (2) $\{y_{t},\ t\geq 0\}$ are strictly increasing; or (3) $\{y_{t},\ t\geq 0\}$ first strictly increase and then become constant. We now prove Theorem 2.7 (a) ($x_{0}<\bar{x}$) by contradiction, showing (Lemma A.3) that structures (1) and (3) are not optimal, so the optimal policy must follow (2). We complete the proof by showing that the repayment amounts go to the limit $\bar{x}$.
Lemma A.3
When $x_{0}<\bar{x}$, both a policy with constant repayment amounts and a policy with strictly increasing repayment amounts followed by constant repayment amounts are sub-optimal.
Proof. The proof is by contradiction. Under both structures (1) and (3), there exists a period $\tau$ where the policy stabilizes with $y_{\tau}=x_{\tau}$. For structure (1), this may start at $\tau=0$ (when $y_{0}=x_{0}$) or at $\tau=1$ (when $y_{0}\neq x_{0}$ and $y_{1}=y_{0}=x_{1}$). For structure (3), $\tau>1$. For each of these cases, we construct an alternative policy with a strictly higher NPV.
(i) $\tau=0$. Here, the optimal policy has a constant repayment amount which is equal to $x_{0}$. The lender’s NPV under this policy is $\Pi_{1}(x_{0})=\frac{1-d}{1-\rho}x_{0}$. Consider an alternative policy with $y_{t}=y$ for all $t\geq 0$. The lender’s NPV under the alternative policy is
$\Pi_{2}(y|x_{0})=x_{0}-dy+\rho\frac{1-F(y)}{1-F(x_{0})}\left(\frac{1-d}{1-\rho}y\right),$
where $\Pi_{1}(x_{0})=\Pi_{2}(x_{0}|x_{0})$. The derivative of $\Pi_{2}(y|x_{0})$ with respect to $y$ is given by
$\frac{d\Pi_{2}(y|x_{0})}{dy}=d\frac{1-F(y)}{1-F(x_{0})}\left(\frac{\rho(1-d)}{d(1-\rho)}-\frac{1-F(x_{0})}{1-F(y)}-\frac{\rho(1-d)}{d(1-\rho)}G(y)\right).$
Since $x_{0}<\bar{x}$, evaluating this derivative at $y=x_{0}$ gives
$\frac{d\Pi_{2}(y|x_{0})}{dy}\bigg{|}_{y=x_{0},x_{0}<\bar{x}}=\frac{\rho(1-d)}{d(1-\rho)}-1-\frac{\rho(1-d)}{d(1-\rho)}G(x_{0})>0,$
so the assumed optimal policy is strictly improved, a contradiction.
(ii) $\tau>0$ (this obviously covers both $\tau=1$ for policy structure (1) and $\tau>1$ for policy structure (3)). We distinguish between two cases: $x_{\tau}<\bar{x}$ and $x_{\tau}\geq\bar{x}$.
Case 1: $x_{\tau}<\bar{x}$. The alternative policy directly generalizes the one for $\tau=0$ in that the repayment amount changes to $y$ starting from period $\tau$ (rather than 00): $y_{t}=y$ for all $t\geq\tau$. We compare the lender’s NPV from period $\tau$ onwards under the original policy, $\Pi_{1}(x_{\tau})$, to that under the alternative policy, $\Pi_{2}(y|x_{\tau})$. Since in this case $x_{\tau}<\bar{x}$, the preceding proof (for $\tau=0$) carries over to this case starting from $\tau$, with the derivative of $\Pi_{2}(y|x_{\tau})$ with respect to $y$ being positive, which shows that the alternative policy strictly improves on the assumed optimal policy, a contradiction.
Case 2. $x_{\tau}\geq\bar{x}$. Under the assumed optimal policy, $x_{\tau-1}<y_{\tau-1}=x_{\tau}=y_{\tau}=x_{\tau+1}=y_{\tau+1}=\dots$. We introduce an alternative policy with repayment amount $y$ for period $\tau-1$ onwards and show that setting $y<y_{\tau-1}$ strictly improve the lender’s NPV. Using similar notation to Case 1 but looking at period $\tau-1$ and all subsequent periods, the lender’s NPV with a repayment amount $y$ is:
${\Pi_{2}(y|x_{\tau-1})=x_{\tau-1}-dy+\rho\frac{1-F(y)}{1-F(x_{\tau}-1)}\left(\frac{1-d}{1-\rho}y\right),}$ (27)
and its derivative with respect to $y$, evaluated at $y\geq\bar{x}$ is
$\frac{d\Pi_{2}(y|x_{\tau-1})}{dy}\bigg{|}_{y\geq\bar{x},y>x_{\tau-1}}=d\frac{1-F(y)}{1-F(x_{\tau-1})}\left(\frac{\rho(1-d)}{d(1-\rho)}-\frac{\rho(1-d)}{d(1-\rho)}G(y)-\frac{1-F(x_{\tau-1})}{1-F(y)}\right)<0,$
which contradicts the optimality of the assumed optimal policy. $\square$
Lemmas A.2 and A.3 together imply that when $x_{0}<\bar{x}$, the optimal policy $\{y_{t},t\geq 0\}$ is strictly increasing. Since the optimal $\{y_{t},t\geq 0\}$ are bounded (by Lemma A.1) and increasing, they have a finite limit. We prove by contradiction that that limit is $\bar{x}$. Suppose that limit is $\hat{x}\neq\bar{x}$. We distinguish between the two cases $\hat{x}<\bar{x}$ and $\hat{x}>\bar{x}$.
Case 1: $\hat{x}<\bar{x}$. Let $\Pi_{t}$ denote the lender’s NPV from period $t$ onwards. Given the limit $\hat{x}$, for any $\epsilon>0$, there exists an integer $M(\epsilon)$ such that
${\hat{x}-x_{t}<\epsilon\quad\textup{and}\quad\hat{x}-y_{t}<\epsilon\quad\textup{for all}\ t\geq M(\epsilon),}$ (28)
and
${|\Pi_{M(\epsilon)}-\frac{1-d}{1-\rho}\hat{x}|<\epsilon,}$ (29)
where $\frac{1-d}{1-\rho}\hat{x}$ is the lender’s NPV from period $M(\epsilon)$ onwards if we set $y_{t}=\hat{x}$ for all $t\geq M(\epsilon)$.
Now consider an alternative policy that sets $y_{t}=y^{\prime}$ for $t\geq M(\epsilon)$. The lender’s NPV from period $M(\epsilon)$ onwards under this policy is given by $\Pi_{2}(y^{\prime}|x_{M(\epsilon)})$, as detailed in equation (27), and $\Pi_{2}(x_{M(\epsilon)}|x_{M(\epsilon)})=\frac{1-d}{1-\rho}x_{M(\epsilon)}$. Moreover, evaluated at $y^{\prime}=x_{M(\epsilon)}$, $\frac{d\Pi_{2}(y^{\prime}|x_{M(\epsilon)})}{dy}>0$, so there exist $\delta,\xi>0$, such that for $y^{\prime}=x_{M(\epsilon)}+\delta$, $\Pi_{2}(y^{\prime}|x_{M(\epsilon)})-\frac{1-d}{1-\rho}x_{M(\epsilon)}>\xi$. Let $\epsilon=\min\{\frac{1}{3}\delta,\frac{1}{2}\frac{1-\rho}{1-d}\xi\}$. Then:
(1) $y^{\prime}=x_{M(\epsilon)}+\delta>x_{M(\epsilon)}+2\epsilon>\hat{x}+\epsilon$, where the first inequality follows since $\epsilon\leq\frac{1}{3}\delta$, and the second inequality follows from (28).
(2) $\Pi_{2}(y^{\prime}|x_{M(\epsilon)})>\frac{1-d}{1-\rho}x_{M(\epsilon)}+\xi>\frac{1-d}{1-\rho}(\hat{x}-\epsilon)+\xi\geq\frac{1-d}{1-\rho}\hat{x}+\frac{1}{2}\xi>\Pi_{M(\epsilon)}-\epsilon+\frac{1}{2}\xi\geq\Pi_{M(\epsilon)}$, where the second inequality follow from equation (28), the third inequality follows from $\epsilon\leq\frac{1}{2}\frac{1-\rho}{1-d}\xi$, the fourth inequality follows from (29), and the last inequality follows from $\epsilon\leq\frac{1}{2}\frac{1-\rho}{1-d}\xi<\frac{1}{2}\xi$, as $\rho>d$. This is a contradiction as the alternative policy improves the lender’s NPV.
Case 2: $\hat{x}>\bar{x}$. We first introduce the following lemma, which shows that whenever $x_{t}\geq\bar{x}$, the optimal policy must be constant from period $t$ onwards.
Lemma A.4
If $x_{t}\geq\bar{x}$ for some $t\geq 0$, the optimal policy must set $y_{k}=x_{t}$ for all $k\geq t$.
Proof. We directly verify the that proposed policy satisfies the Bellman equation, where
${J(x_{t})=\frac{1-d}{1-\rho}x_{t}=\max_{y_{t}\in[x_{t},u)}\ x_{t}-dy_{t}+\rho\frac{1-F(y_{t})}{1-F(x_{t})}J(y_{t}).}$ (30)
Substituting of $J(y_{t})=\frac{1-d}{1-\rho}y_{t}$ into equation (30), we obtain the maximization problem
${\max_{y_{t}\in[x_{t},u)}\ x_{t}-dy_{t}+\rho\frac{1-F(y_{t})}{1-F(x_{t})}\frac{1-d}{1-\rho}y_{t}.}$ (31)
The derivative of the objective function in (31) with respect to $y_{t}$ is $d\frac{1-F(y)}{1-F(x_{t})}\left(\frac{\rho(1-d)}{d(1-\rho)}-\frac{1-F(x_{t})}{1-F(y)}-\frac{\rho(1-d)}{d(1-\rho)}G(y)\right)$, which is negative for any $x_{t}\geq\bar{x}$ and $y_{t}>x_{t}$. It follows that $y_{t}=x_{t}$ solves the Bellman equation, and equation (30) holds. The optimal policy is thus to set $y_{k}=x_{t}$ for all $k\geq t$, which yields $J(x_{t})=\frac{1-d}{1-\rho}x_{t}$. $\square$
Note that Lemma A.4 holds in general and need not assume $x_{0}<\bar{x}$. However, we are assuming here that $x_{0}<\bar{x}$ and by Lemma A.3, a series of constant $y_{t}$ is not optimal, a contradiction.
We have thus proven that the limit of the repayment amounts must be $\bar{x}$, which completes the proof of Theorem 2.7 (a). Lemma A.4 in fact proves Theorem 2.7 (b) by setting $t=0$: It shows that when $x_{0}\geq\bar{x}$, the optimal policy is to set $y_{k}=x_{0}$ for all $k\geq$0. This completes the proof of Theorem 2.7. $\square$
A.5 Proof of Theorem 2.5
In the uniform case, the Bellman equation becomes
${J(x_{t})=\max_{y_{t}}\ x_{t}-dy_{t}+\rho\dfrac{1-y_{t}}{1-x_{t}}J(y_{t}).}$
We’ll show that the proposed linear control satisfies the Bellman equation. Given the existence and uniqueness of the optimal solution (Lemma A.1), this will prove the Theorem.
With a linear control, $J(x)$ is of the form [^1] For notational simplicity, we suppress the dependence on the time period.
${J(x)=\begin{cases}x+\dfrac{ax^{2}+bx+c}{1-x},&\textup{if }x<\bar{x},\\ \dfrac{(1-d)}{(1-\rho)}x,&\textup{if }x\geq\bar{x},\end{cases}}$
where $\bar{x}$ is the unique solution to (4), which is reduced to ${\bar{x}=\frac{(\rho-d)}{2\rho-\rho d-d}}$ in the uniform case. We can now use the following conditions to solve for ${\{a,b,c\}}$ (all for $x\in(0,\bar{x})$ ):
Value Matching:
${J(\bar{x})=\bar{x}+\dfrac{a\bar{x}^{2}+b\bar{x}+c}{1-\bar{x}}=\dfrac{(1-d)}{(1-\rho)}\bar{x};}$ (32)
Smoothness:
${J^{\prime}(\bar{x})=\dfrac{a\bar{x}^{2}+b\bar{x}+c}{(1-\bar{x})^{2}}+\dfrac{2a\bar{x}+b}{1-\bar{x}}+1=\dfrac{(1-d)}{(1-\rho)};}$ (33)
Linear Control:
$\begin{split}\dfrac{\partial J(x)}{\partial y}=&\dfrac{(x+(\rho(1+b)/d-1))-2\rho((1-a)/d)y}{(1-x)/d}=0,\\ y=&mx+n,\quad m=\dfrac{d}{2\rho(1-a)},\quad n=\dfrac{(\rho-d+b\rho)}{2\rho(1-a)},\\ \bar{x}=&m\bar{x}+n;\end{split}$ (34)
Consistency of Bellman equation at $x=0$:
${J(0)=c=-dn+\rho\left(1-n\right)\left(n+\dfrac{an^{2}+bn+c}{1-n}\right).}$ (35)
With some algebra, we find two sets ${\{a_{1},b_{1},c_{1}\}}$ and ${\{a_{2},b_{2},c_{2}\}}$ that satisfy (32)-(35):
${\begin{array}[]{cl}a_{1}=&\dfrac{1}{2}\left(1+d\sqrt{\dfrac{\rho-d^{2}}{\rho d^{2}}}\right),\\ b_{1}=&\dfrac{1}{2\rho-\rho d-d}\left((\rho-d)\left(-d\sqrt{\dfrac{\rho-d^{2}}{\rho d^{2}}}+d-1\right)\right),\\ c_{1}=&\dfrac{1}{2(1-\rho)(2\rho-d-\rho d)^{2}}\left((\rho-d)^{2}\left(1-2d+\rho+(1-\rho)\sqrt{1-\dfrac{d^{2}}{\rho}}\right)\right);\\ a_{2}=&\dfrac{1}{2}\left(1-d\sqrt{\dfrac{\rho-d^{2}}{\rho d^{2}}}\right),\\ b_{2}=&\dfrac{1}{2\rho-\rho d-d}\left((\rho-d)\left(d\sqrt{\dfrac{\rho-d^{2}}{\rho d^{2}}}+d-1\right)\right),\\ c_{2}=&\dfrac{1}{2(1-\rho)(2\rho-d-\rho d)^{2}}\left((\rho-d)^{2}\left(1-2d+\rho-(1-\rho)\sqrt{1-\dfrac{d^{2}}{\rho}}\right)\right).\end{array}}$
By the linear control requirement, $(1-m)\bar{x}=n$, and since $\bar{x}$ and $n$ are positive, we must have $m<1$. The first solution does not satisfy this requirement since $\rho>d$ implies $m(a_{1})=\dfrac{d}{2\rho(1-a_{1})}>1$.
We complete the proof by showing that $\{a_{2},b_{2},c_{2}\}$ satisfy the Bellman equation. For $x\in(0,\bar{x})$, ${J(x_{t})}$ should satisfy
${J(x_{t})=x_{t}+\dfrac{ax_{t}^{2}+bx_{t}+c}{1-x_{t}}=\max_{y_{t}\in[x_{t},u)}x_{t}-dy_{t}+\rho\dfrac{u-y_{t}}{u-x_{t}}J(y_{t})=\max_{y_{t}\in[x_{t},u)}x_{t}-dy_{t}+\rho\dfrac{u-y_{t}}{u-x_{t}}\left(y_{t}+\dfrac{ay_{t}^{2}+by_{t}+c}{1-y_{t}}\right).}$ (36)
Plugging in $\{a_{2},b_{2},c_{2}\}$, equation (36) holds. $\square$
A.6 Proof of Proposition 2.6
We first prove that if in period $k^{*}$, $\Pi_{k^{*}}\geq 0$, then $\Pi_{k+1}>\Pi_{k}\geq 0$ for all $k\geq k^{*}$. This implies that unprofitable consumers default in periods $k=1,2,\ldots,k^{*}-1$, profitable consumers default in periods $k\geq k^{*}$, and creditworthy consumers never default; and that there exists a cutoff value $\theta^{*}(\rho,d)=y_{k^{*}}$ which separates unprofitable from profitable consumers.
As $\Pi_{k}=\sum^{k-1}_{i=0}\rho^{i}(\rho-d)y_{i}-\rho^{k}dy_{k}$, $\Pi_{k+1}-\Pi_{k}=y_{k}-dy_{k+1}$ and $y_{k}=mx_{k}+n$ for the uniform case. With some algebra, $\frac{1-d}{d}>\frac{m^{k+1}}{\sum^{k}_{i=0}m^{i}}$ implies that $y_{k}-dy_{k+1}>0$, which further implies $\Pi_{k+1}>\Pi_{k}$. As ${\frac{m^{k+1}}{\sum^{k}_{i=0}m^{i}}=\frac{1-m}{m^{-(k+1)}-1}}$ is decreasing in $k$, if $\frac{1-d}{d}>\frac{1-m}{m^{-k^{*}}-1}$ in period $k^{*}$, then $\frac{1-d}{d}>\frac{1-m}{m^{-k+1}-1}$ and therefore $\Pi_{k+1}>\Pi_{k}$ for all $k\geq k^{*}$.
[…]
Since $k^{*}$ is the smallest $k^{\prime}$ such that $\Pi_{k^{\prime}}\geq 0$, $\theta^{*}(\rho,d)=y_{k^{*}}$. Consumers with $\theta\in[0,\theta^{*}(\rho,d))$ are unprofitable, and those with $\theta\in[\theta^{*}(\rho,d),\theta^{\infty}(\rho,d))$ are profitable. Consumers with $\theta\in[\theta^{\infty}(\rho,d),1)$ are credit-worthy with $\theta^{\infty}(\rho,d)=\bar{x}$.
To complete the proof, we show that the three segments are nonempty. The unprofitable customers are from the interval $\theta\in[0,y_{0})$, which is nonempty since $y_{0}=n>0$. Also, since there always exists a $k^{*}$ such that $\Pi_{k^{*}}\geq 0$ and $y_{k}^{*}<\bar{x}$, the set of profitable but non-creditworthy consumers is also nonempty. Finally, by Lemma 2.3, $\bar{x}<u$, so the set of creditworthy consumers is notempty. $\square$
A.7 Proof of Lemma 3.1
Similar to the proof of Lemma 2.2, if the lender sets $y_{t}<x_{t}$ with discount rate $d_{t}$, its expected NPV will be ${-s(d_{t})d_{t}y_{t}+\rho s(d_{t})\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})[y_{t}+J(x_{t})]=s(d_{t})((\rho-d_{t})y_{t}+\rho J(x_{t})),}$ since $\operatorname{\mathbb{P}}(\theta\geq y_{t}|\theta\geq x_{t})=1$. The NPV with $\tilde{y}_{t}=y_{t-1}=x_{t}$ under the same discount rate $d_{t}$ is ${s(d_{t})\left(\left(\rho-d_{t}\right)x_{t}+\rho J(x_{t})\right)}$. The difference between the two NPV values is $s(d_{t})\left(\rho-d_{t}\right)(x_{t}-y_{t})>0$. Since this inequality holds for all $t$, it’s never optimal to set $y_{t}<x_{t}$. $\square$
A.8 Proof of Lemma 3.2
As $\rho>d$, $-\frac{(1-\rho s(d^{*}))}{(1-s(d^{*})d^{*})}\frac{d^{*}}{\rho}+1\in(0,1)$ for any $d^{*}$, so there is a unique $\bar{x}$ solving equation (13). As for $d^{*}$, we rearrange (14) as $(\rho-d)s^{\prime}(d)+\rho s(d)^{2}-s(d)=0$. The left hand side of this equation is monotone decreasing (with the derivative $(\rho-d)s^{\prime\prime}(d)+2s^{\prime}(d)(\rho s(d)-1)<0$); $\lim_{d\to 0}(\rho-d)s^{\prime}(d)+\rho s(d)^{2}-s(d)>0$; and $\lim_{d\to\rho}(\rho-d)s^{\prime}(d)+\rho s(d)^{2}-s(d)<0$. It follows that (14) has a unique solution.
We complete the proof by showing that with constant demand elasticity ($s(d)=d^{\alpha}$), $d^{*}$ is an increasing function of $\alpha$. With a constant elasticity, (15) is equivalent to $(\rho-d)\alpha d^{\alpha-1}+\rho d^{2\alpha}-d^{\alpha}=0$. By the implicit function theorem,
${\frac{\partial d^{*}}{\partial\alpha}=\frac{\rho(d^{*})^{\alpha+1}log(d^{*})+\rho-d^{*}}{(\alpha+1)(1-\rho(d^{*})^{\alpha})}.}$
To prove that $\frac{\partial d^{*}}{\partial\alpha}>0$. it’s sufficient to show that the numerator ${\rho(d^{*})^{\alpha+1}log(d^{*})+\rho-d^{*}}$ is positive. By (14), $\rho=\frac{\alpha+1}{\alpha+(d^{*})^{\alpha+1}}d^{*}$, and the numerator equals ${\frac{d^{*}}{\alpha+(d^{*})^{\alpha+1}}\left(1-(d^{*})^{\alpha+1}(1-log(d^{*})(\alpha+1))\right).}$ Let $\delta(d)=d^{\alpha+1}(1-log(d)(\alpha+1))$, so $\delta^{\prime}(d)=-(\alpha+1)^{2}d^{\alpha}log(d)$, which is positive as $d\in(0,\rho)$. Further, $\delta(1)=1$, $01-(d^{*})^{\alpha+1}(1-log(d^{*})(\alpha+1))=\delta(1)-\delta(d^{*})>$ 0. The numerator is thus given by $\frac{d^{*}}{\alpha+(d^{*})^{\alpha+1}}(\delta(1)-\delta(d))>0$. $\square$
A.9 Proof of Theorem 3.3
Theorem 3.3 is generalized in the analysis of the hybrid architecture. To keep the paper compact, we relegated the analysis of the hybrid case to the Appendix. Here, we state and prove the Theorem for the hybrid case (Theorem A.5), with Theorem 3.3 being a special case. As outlined below, the optimal policy and the proof are similar to those of Theorem 2.7 in Appendix A.4.
Theorem A.5
For the hybrid architecture with a monotone decreasing demand elasticity, there exists a unique optimal policy $\BFysuperscript\BFd(\BFy^{*},\BFd^{*})$ with the following structure:
  • (a) If $x_{0}<\bar{x}$, the lender keeps increasing the repayment amounts $y_{t}$ in each period $t$ as long as the consumer repays her loans, and stops lending if the consumer defaults. The infinite sequence of (potential) repayment amounts $\{y_{t},t=0,1,2,\ldots\}$ is strictly increasing to a limit $\bar{x}$ defined by equation (13).
  • (b) If $x_{0}\geq\bar{x}$, the optimal policy is to set the repayment amount at $y_{t}=x_{0}$ in each period with a fixed discount rate $d^{*}$.
The proof is similar to the proof of Theorem 2.7 (Appendix A.4). We first establish the existence and uniqueness of the solution to the Bellman equation as in Lemma A.1: since both controls are bounded and the discount factor $\rho\in(0,1)$, there exists a unique solution to the Bellman equation. We next show that if the repayment amounts $y_{t}$ satisfy $y_{t}=y_{t-1}=x_{t}$, then the optimal policy sets $y_{k}=x_{k}$ for all subsequent periods with $d_{k}=d^{*}$. If $y^{*}_{t}=x_{t}$, then $x_{t}$ maximizes the Bellman equation in period $t$: ${J(x_{t})\ =\ \max_{y_{t}\in[x_{t},u),d_{t}\in(0,\rho)}\ \ {x_{t}-s(d_{t})d_{t}y_{t}+\rho s(d_{t})\dfrac{1-F(y_{t})}{1-F(x_{t})}J(y_{t})}.}$ Now, the Bellman equation for period $t+1$ is the same as for period $t$. Hence, $y=x_{t}$ also maximizes the Bellman equation for period $t+1$, so $y^{*}_{t+1}=y^{*}_{t}=x_{t}$. Further, when the repayment amount stabilizes at $x_{t}$, the lender’s expected NPV is given by $\frac{1-s(d_{t})d_{t}}{1-s(d_{t})\rho}x_{t}$, with $d^{*}$ being the unique maximizer. The result now follows by forward induction.
It follows that the optimal repayment amounts have to follow one of the same three structures ((1)-(3)) we identified in Appendix A.4. The following Lemma, analogous to Lemma A.3 in Appendix A.4, shows that the optimal repayment amounts must be strictly increasing. We then show that the limit of the repayment amounts must be $\bar{x}$.
Lemma A.6
When $x_{0}<\bar{x}$, under an endogenous interest rate and decreasing demand elasticity, both a policy with constant repayment amounts and a policy with strictly increasing repayment amounts followed by constant repayment amounts are sub-optimal.
Proof. The proof is by contradiction. Under both structures, there exists a period $\tau$ where the policy stabilizes with $y_{\tau}=x_{\tau}$. We construct alternative policies with strictly higher NPVs for $\tau=0$ and $\tau>0$.
(i) $\tau=0$. Here, the optimal policy has a constant repayment amount $y_{t}=x_{0}$, with $d_{t}=d^{*}$, and the lender’s NPV is ${\Pi_{1}(d^{*},x_{0})=\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}x_{0}=x_{0}-s(d^{*})d^{*}x+\rho s(d^{*})\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}x_{0}}$. We now consider an alternative policy which sets the repayment amounts at $y$ for all $t\geq 0$ with the same discount rate $d^{*}$. The lender’s NPV under the alternative policy is
${\Pi_{2}(d^{*},y|x_{0})=x_{0}-s(d^{*})d^{*}y+\rho s(d^{*})\frac{1-F(y)}{1-F(x_{t})}\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}y,}$
where $\Pi_{1}(d^{*},x_{0})=\Pi_{2}(d^{*},x_{0}|x_{0})$. The derivative of $\Pi_{2}(d^{*},y|x_{0})$ with respect to $y$ is given by
${\frac{\partial\Pi_{2}(d^{*},y|x_{0})}{\partial y}=-s(d^{*})d^{*}+\frac{\rho s(d^{*})(1-s(d^{*})d^{*})}{1-s(d^{*})\rho}\left(\frac{1-F(y)}{1-F(x_{t})}-\frac{f(y)}{1-F(x_{t})}y\right).}$
If $x_{0}<\bar{x}$, evaluating this derivative at $y=x_{0}$ gives,
${\frac{d\Pi_{2}(d^{*},y|x_{0})}{dy}\bigg{|}_{y=x_{0},x_{0}<\bar{x}}=\frac{\rho s(d^{*})(1-s(d^{*})d^{*})}{1-s(d^{*})\rho}\left(1-\frac{d^{*}(1-s(d^{*})\rho)}{\rho(1-s(d^{*})d^{*})}-G(x)\right)>0.}$
This implies that when $x_{0}<\bar{x}$, the assumed optimal policy is strictly improved by deviating to a larger constant value $y>x_{0}$ for all $t\geq 0$ with the same discount rate $d^{*}$, a contradiction.
(ii) $\tau>0$. We distinguish between two cases : $x_{\tau}<\bar{x}$ and $x_{\tau}\geq\bar{x}$.
Case 1. $x_{\tau}<\bar{x}$. Here, the alternative policy directly generalizes the one for $\tau=0$ in that the repayment amount changes to $y$ starting from period $\tau$ rather than 00, and the above proof follows through.
Case 2. $x_{\tau}\geq\bar{x}$. Looking at period $\tau-1$ and all subsequent periods, under the assumed policy where $x_{\tau-1}<y_{\tau-1}=x_{\tau}=y_{\tau}=x_{\tau+1}=y_{\tau+1}=\dots$, the lender’s NPV with repayment amount $y_{\tau-1}=y$ and discount rate $d_{\tau}$ is
${\Pi_{2}(d_{\tau-1},y|x_{\tau-1})=x_{\tau-1}-s(d_{\tau-1})d_{\tau-1}y_{\tau-1}+\rho s(d_{\tau-1})\frac{1-F(y)}{1-F(x_{\tau-1})}\left(\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}y\right).}$ (37)
The discount rate $d_{\tau-1}$ that maximizes the lender’s NPV satisfies
${\frac{1-F(x_{\tau-1})}{1-F(y)}=\frac{s(d^{*})+s^{\prime}(d^{*})d^{*}}{\rho s^{\prime}(d^{*})}\frac{\rho s^{\prime}(d_{2})}{s(d_{\tau-1})+s^{\prime}(d_{\tau-1})d_{\tau-1}}.}$ (38)
Further, $d_{\tau-1}<d^{*}$. To see this, let $\eta(d)=\frac{s(d)+s^{\prime}(d)d}{\rho s^{\prime}(d)}$. Equation (38) becomes $\eta(d_{\tau-1})=\frac{1-F(x_{\tau-1}+\epsilon)}{1-F(x_{\tau-1})}\eta(d^{*})$. Moreover, $\eta(d)$ is monotone increasing as $\eta^{\prime}(d)=\frac{2\rho s^{\prime}(d)-\rho s^{\prime\prime}(d)s(d)}{(\rho s^{\prime}(d))^{2}}>0$. Hence $d_{\tau-1}<d^{*}$.
We now take the derivative of $\Pi_{2}$ with respect to $y$. By the evenlope theorem, $\frac{d\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{dy}=\frac{\partial\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{\partial y}+\frac{\partial\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{\partial d_{\tau-1}}\frac{\partial d_{\tau-1}}{\partial y}$, and $\frac{\partial\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{\partial d_{\tau-1}}=0$. The derivative $\frac{d\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{dy}$ is thus given by
${\frac{d\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{dy}=\rho s(d_{\tau-1})\frac{1-F(y)}{1-F(x_{\tau-1})}\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}\left[1-\frac{1-F(x_{\tau-1})}{1-F(y)}\frac{1-s(d^{*})\rho}{1-s(d^{*})d^{*}}\frac{d_{\tau-1}}{\rho}-G(y)\right].}$ (39)
Plugging (38) into the derivative (39), we have
${\frac{d\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{dy}=\rho s(d_{2})\frac{1-F(y)}{1-F(x_{\tau-1})}\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}\left[1-\left(\frac{s(d^{*})+s^{\prime}(d^{*})d^{*}}{s^{\prime}(d^{*})}\frac{s^{\prime}(d_{\tau-1})}{s(d_{\tau-1})+s^{\prime}(d_{\tau-1})d_{\tau-1}}\frac{d_{\tau-1}}{d^{*}}\right)\frac{1-s(d^{*})\rho}{1-s(d^{*})d}\frac{d^{*}}{\rho}-G(y)\right].}$
However, evaluating at $y\geq\bar{x}$,
${\frac{d\Pi_{2}(d_{\tau-1},y|x_{\tau-1})}{dy}\bigg{|}_{y\geq\bar{x}}<0.}$ (40)
This is because as the elasticity is decreasing and $d_{\tau-1}<d^{*}$, ${\frac{s(d^{*})+s^{\prime}(d^{*})d^{*}}{s^{\prime}(d^{*})}\frac{s^{\prime}(d_{\tau-1})}{s(d_{\tau-1})+s^{\prime}(d_{\tau-1})d_{\tau-1}}\frac{d_{\tau-1}}{d^{*}}>1}$. Following the definition of $G(\bar{x})$ (equation (14), Lemma 3.2), equation (40) holds. This implies that the assumed optimal policy can be strictly improved (by decreasing the repayment amount $y_{\tau-1}$), a contradiction. $\square$
The previous results imply that when $x_{0}<\bar{x}$, the optimal repayment amounts $y_{t}$ are strictly increasing. We next show that the limit of the repayment amount sequence is $\bar{x}$. Suppose by contradiction that the limit is $\hat{x}\neq\bar{x}$. We then distinguish between two cases, $\hat{x}<\bar{x}$ and $\hat{x}>\bar{x}$.
Case 1: $\hat{x}<\bar{x}$. The proof follows that of Appendix A.4, replacing $|\Pi_{M(\epsilon)}-\frac{1-d}{1-\rho}\hat{x}|<\epsilon|$ by $|\Pi_{M(\epsilon)}-\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}\hat{x}|<\epsilon|$, where $\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}\hat{x}$ is the lender’s NPV from period $M(\epsilon)$ onwards with $y_{t}=\hat{x}$ and $d_{t}=d^{*}$. Further, we replace $\Pi_{2}(y^{\prime}|x_{M(\epsilon)})$ with $\Pi_{2}(d^{*},y^{\prime}|x_{M(\epsilon)})$ (equation (37), which is the counterpart of equation (27) in Appendix A.4).
Case 2: $\hat{x}\geq\bar{x}$. We first introduce the following Lemma, which states that whenever $x_{t}\geq\bar{x}$, the optimal policy becomes constant with repayment amount and discount rate $(x_{t},d^{*})$ from period $t$ onwards.
Lemma A.7
If $x_{t}\geq\bar{x}$ for some $t\geq 0$, then the optimal policy is to set $(y_{k},d_{k})=(x_{t},d^{*})$ for all $k\geq t$.
Proof. We directly verify the optimality of the Bellman equation,
${J(x_{t})=\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}x_{t}=\max_{y_{t}\in[x_{t},u),d_{t}\in(0,\rho)}\ x_{t}-s(d)dy_{t}+\rho s(d)\frac{1-F(y_{t})}{1-F(x_{t})}J(y_{t}).}$ (41)
Substituting $J(y_{t})=\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}y_{t}$, the right hand side of (41) becomes
${\max_{y_{t}\in[x_{t},u),d_{t}\in(0,\rho)}\ x_{t}-ds(d)y_{t}+\rho s(d)\frac{1-F(y_{t})}{1-F(x_{t})}\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}y_{t}.}$ (42)
The optimal $d_{t}$ under any $y_{t}$ should satisfy (38) (by changing $d_{\tau-1}$ in equation (38) to $d_{t}$). By (40), $\frac{d\Pi_{2}(d_{t},y|x_{t})}{dy}\bigg{|}_{y>x_{t}\geq\bar{x}}<0$. Any $y>x_{t}$ is not an optimal solution of the problem (42), hence $y_{t}=x_{t}$ and $d_{t}=d^{*}$ is the unique pair that solves the Bellman equation when $x_{t}\geq\bar{x}$, and equation (41) holds. The optimal policy is to set $y_{k}=x_{t}$, $d_{k}=d^{*}$ for all $k\geq t$, and under this policy, $J(x_{t})=\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}x_{t}$. $\square$
Lemma A.7 shows that under our assumptions, the repayment amount becomes constant, but Lemma A.6 shows that when $x_{0}<\bar{x}$, this solution cannot be optimal, a contradiction.
We have thus proved that the limit of the repayment amounts is $\bar{x}$, which completes the proof of Theorem A.5 (a). Lemma A.7 also proves Theorem A.5 (b) by setting $t=0$. This completes the proof of Theorem A.5, hence also Theorem 3.3. $\square$
A.10 Proof of Theorem 3.4
We prove Theorem 3.6 (Theorem 3.4 is a special case) by directly verifying that the proposed solution satisfies the Bellman equation. For notational simplicity, we omit here the dependence on $t$ and just write $x$ instead of $x_{t}$. The results hold for all $t$. We also define $\beta=\frac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}$, so equation (14) becomes $\beta=\frac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}=\frac{1+\alpha}{\rho\alpha}d^{*}$.
We consider separately $x<\bar{x}$ and $x\geq\bar{x}$.
Case 1: $x<\bar{x}$. We verify that the following equation holds,
${\begin{array}[]{cl}J(x)|_{x<\bar{x}}&=x-(d_{x})^{\alpha+1}\bar{x}+\rho(d_{x})^{\alpha}\frac{1-F(\bar{x})}{1-F(x)}\frac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}\bar{x}\\ &=\max\left\{\max_{d,y\in[x,\bar{x})}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right),\max_{d,y\in[\bar{x},u)}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\beta y\right\},\end{array}}$
where
${d_{x}=\frac{1-F(\bar{x})}{1-F(x)}d^{*}\quad\textup{and }\quad d_{y}=\frac{1-F(\bar{x})}{1-F(y)}d^{*}.}$
This involves two cases: evaluating $\max_{d,y\in[x,\bar{x})}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right)$ for $y\in[x,\bar{x})$ (Case 1.1 below), and evaluating $\max_{d,y\in[\bar{x},u)}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\beta y$ for $y\in[\bar{x},u)$ (Case 1.2 below).
Case 1.1: $y\in[x,\bar{x})$. We prove that
${\max_{d,y\in[x,\bar{x})}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right)=x-d_{x}^{\alpha+1}\bar{x}+\rho d_{x}^{\alpha}\frac{1-F(\bar{x})}{1-F(x)}\frac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}\bar{x}.}$ (43)
Consider the optimization problem
${\max_{d,y\in[x,\bar{x})}\Pi_{1}(d,y)\ :=x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right).}$
We first optimize over $d$. The optimal $d$ is always in the interior of $(0,\rho)$ and it satisfies the first order condition ${-(\alpha+1)d^{\alpha}y+\rho\alpha d^{\alpha-1}\frac{1-F(y)}{1-F(x)}\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right)=0.}$ With some algebra,
${d=\frac{\rho\alpha}{\alpha+1}\frac{1-F(y)}{1-F(x)}\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right)\frac{1}{y}=\frac{\rho\alpha}{\alpha+1}\frac{1-F(y)}{1-F(x)}\frac{C(y)}{y},}$ (44)
where
${C(y)=\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right).}$
We next optimize over $y$, taking the derivative of $\Pi_{1}$ with respect to $y$: ${\frac{d\Pi_{1}}{dy}=\frac{\partial\Pi_{1}}{\partial y}+\frac{\partial\Pi_{1}}{\partial d}\frac{\partial d}{\partial y}.}$ By the envelope theorem, $\frac{\partial\Pi_{1}}{\partial d}=0$, and
$\frac{d\Pi_{1}}{dy}=\frac{\partial\Pi_{1}}{\partial y}$ $=-d^{\alpha+1}+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\frac{dC(y)}{dy}-\rho d^{\alpha}\frac{f(y)}{1-F(x)}C(y).$ (45)
Substituting (44) into (45), we obtain
${\frac{d\Pi_{1}}{dy}=\frac{\rho d^{\alpha}}{y}\frac{1-F(y)}{1-F(x)}\left(-\frac{\alpha}{\alpha+1}C(y)+y\frac{dC(y)}{dy}-G(y)C(y)\right).}$
By the definition of $d_{y}$, $\frac{\partial C(y)}{\partial d_{y}}=0$, which implies,
${\frac{dC(y)}{dy}=\frac{\partial C(y)}{\partial y}+\frac{\partial C(y)}{\partial d_{y}}\frac{\partial d_{y}}{\partial y}=\frac{\partial C(y)}{\partial y}=1+\rho d_{y}^{\alpha}\frac{1-F(\bar{x})}{1-F(y)}\frac{f(y)}{1-F(y)}\beta\bar{x}.}$
Further,
$\frac{d\Pi_{1}}{dy}$ $=\frac{d^{\alpha}\rho}{y}\frac{1-F(y)}{1-F(x)}\left(-\frac{\alpha}{\alpha+1}C(y)+y\left(1+\rho d_{y}^{\alpha}\frac{1-F(\bar{x})}{1-F(y)}\frac{f(y)}{1-F(y)}\beta\bar{x}\right)-G(y)C(y)\right)$
$=\frac{d^{\alpha}\rho}{y}\frac{1-F(y)}{1-F(x)}\left(-\frac{\alpha}{\alpha+1}C(y)+y+G(y)\left(\rho d_{y}^{\alpha}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right)-G(y)C(y)\right)$
$=\frac{d^{\alpha}\rho}{y}\frac{1-F(y)}{1-F(x)}\left(-\left(\frac{\alpha}{\alpha+1}+G(y)\right)\left(y-d^{\alpha+1}_{y}\bar{x}\right)+y-\frac{\alpha}{\alpha+1}\left(\rho d_{y}^{\alpha}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right)\right).$
By definition, $d_{y}=\frac{1-F(\bar{x})}{1-F(y)}d^{*}=\frac{\alpha\rho}{\alpha+1}\frac{1-F(\bar{x})}{1-F(y)}\beta$, and
$\frac{d\Pi_{1}}{dy}$ $=\frac{d^{\alpha}\rho}{y}\frac{1-F(y)}{1-F(x)}\left(-\left(\frac{\alpha}{\alpha+1}+G(y)\right)\left(y-d^{\alpha+1}_{y}\bar{x}\right)+y-d^{\alpha+1}_{y}\bar{x}\right)$
$=\frac{d^{\alpha}\rho}{y}\frac{1-F(y)}{1-F(x)}\left(1-\left(\frac{\alpha}{\alpha+1}+G(y)\right)\right)\left(y-d^{\alpha+1}_{y}\bar{x}\right).$
Since $y<\bar{x}$, ${G(y)<-\frac{\alpha}{\alpha+1}+1,}$ hence $\left(1-\left(\frac{\alpha}{\alpha+1}+G(y)\right)\right)>0$.
Let $q(y)=y-d^{\alpha+1}_{y}\bar{x}$, then
${\frac{d\Pi_{1}}{dy}=\frac{d^{\alpha}\rho}{y}\frac{1-F(y)}{1-F(x)}\left(1-\left(\frac{\alpha}{\alpha+1}+G(y)\right)\right)q(y)}$
for $y\in[x,\bar{x})$. At $y=\bar{x}$, $q(y)|_{y=\bar{x}}=y-d^{\alpha+1}_{y}\bar{x}=(1-d^{\alpha+1}_{\bar{x}})\bar{x}>0$. Moreover, $q^{\prime}(y)=1-d^{\alpha}_{y}\alpha\rho\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}$ is monotone decreasing in $y$. Together with the fact that $q(0)<0$, $q(\bar{x})>0$ and $q^{\prime}(y)$ is monotone decreasing, $q(y)$ crosses the $x-$axis only once for $x\in[0,\bar{x})$. This implies the single-crossing property for $\frac{d\Pi_{1}}{dy}(y)$: there exists a $y^{\prime}\in(0,\bar{x})$ such that for $y<y^{\prime}$, $\frac{d\Pi_{1}}{dy}<0$, and for $y>y^{\prime}$, $\frac{d\Pi_{1}}{dy}>0$
We further prove that $\frac{d\Pi_{1}}{dy}|_{y=x}>0$ for all $x<\bar{x}$:
${\frac{d\Pi_{1}}{dy}\bigg{|}_{y=x}=(d^{*})^{\alpha}\rho\beta\left(-\frac{\alpha}{\alpha+1}+1-G(x)\right)>0,\quad\textup{for }x<\bar{x}.}$
Finally, by the single-crossing property of $\frac{d\Pi_{1}}{dy}$, $\frac{d\Pi_{1}}{dy}>0$ for $y\in(x,\bar{x}]$. This implies that $y=\bar{x}$ optimizes $\Pi_{1}$, and equation (43) holds.
Case 1.2: $y\in[\bar{x},u)$. We prove that
${\max_{d,y\in[\bar{x},u)}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\beta y=x-d_{x}^{\alpha+1}\bar{x}+\rho d_{x}^{\alpha}\frac{1-F(\bar{x})}{1-F(x)}\beta\bar{x}.}$ (46)
Consider the optimization problem
${\max_{d,y\in[\bar{x},u)}\ \Pi_{2}(d,y):=x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\beta y.}$ (47)
We first optimize over $d$, obtaining ${d=\frac{\alpha\rho}{\alpha+1}\frac{1-F(y)}{1-F(x)}\beta=\frac{1-F(y)}{1-F(x)}d^{*}.}$
We next optimize over $y$. The derivative of $\Pi_{2}$ with respect to $y$ is ${\frac{d\Pi_{2}}{dy}=\frac{\partial\Pi_{2}}{\partial y}+\frac{\partial\Pi_{2}}{\partial d}\frac{\partial d}{\partial y}.}$ By the envelope theorem, $\frac{\partial\Pi_{2}}{\partial d}=0$, and
${\frac{d\Pi_{2}}{dy}=\frac{\partial\Pi_{2}}{\partial y}=d^{\alpha}\left(-d+\rho\frac{1-F(y)}{1-F(x)}\beta-\rho\beta\frac{f(y)}{1-F(x)}y\right)=d^{\alpha}\rho\beta\frac{1-F(y)}{1-F(x)}\left(-\frac{\alpha}{\alpha+1}+1-G(y)\right).}$
Since $G(y)>1-\frac{d^{*}}{\rho\beta}=1-\frac{\alpha}{\alpha+1}$ for all $y>\bar{x}$ and $G(y)=1-\frac{\alpha}{\alpha+1}$ for $y=\bar{x}$, it follows that ${\frac{d\Pi_{2}}{dy}|_{y>\bar{x}}<0}$ and ${\frac{d\Pi_{2}}{dy}|_{y=\bar{x}}=0.}$ Hence, equation (46) holds, and for $x<\bar{x}$,
$J(x)|_{x<\bar{x}}=$ $x-d_{x}^{\alpha+1}\bar{x}+\rho d_{x}^{\alpha}\frac{1-F(\bar{x})}{1-F(x)}\frac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}\bar{x}$
$=$ $\max\left\{\max_{d,y\in(x,\bar{x})}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\left(y-d^{\alpha+1}_{y}\bar{x}+\rho d^{\alpha}_{y}\frac{1-F(\bar{x})}{1-F(y)}\beta\bar{x}\right),\quad\max_{d,y\in[\bar{x},u)}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\beta y\right\}$
with corresponding ${d_{x}=\frac{1-F(\bar{x})}{1-F(x)}d^{*},\quad y_{x}=\bar{x}.}$
Case 2: $x\geq\bar{x}$.
We verify that the following equation holds:
${J(x)|_{x\geq\bar{x}}=\beta x=\max_{d,y\in[x,u)}\ x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\beta y.}$ (48)
The proof is similar to that of Case 1.2. Consider the optimization problem
${\max_{d,y\in[x,u)}\Pi_{2}(d,y)=x-d^{\alpha+1}y+\rho d^{\alpha}\frac{1-F(y)}{1-F(x)}\beta y.}$ (49)
The optimal $d=\frac{1-F(y)}{1-F(x)}d^{*}$, and the derivative of $\Pi_{2}$ with respect to $y$ is
${\frac{d\Pi_{2}}{dy}=d^{\alpha}\rho\beta\frac{1-F(y)}{1-F(x)}\left(-\frac{\alpha}{\alpha+1}+1-G(y)\right).}$ (50)
For $y>x>\bar{x}$, $\frac{d\Pi_{2}}{dy}|_{y>\bar{x}}<0$. Hence, the optimal $y$ is attained at $y=x$, with $d_{x}=d^{*}$, and equation (48) holds.
We have thus verified the optimality of the proposed solution for Cases 1 and 2, which completes the proof. $\square$
A.11 Proof of Theorem 3.5 and its Hybrid generalization
Theorem 3.5 is generalized in the analysis of the hybrid architecture. As before, we state and prove the Theorem for the hybrid case here.
Theorem A.8
[…]
  • (a) if $x_{0}<\bar{x}$, in period 00 the lender sets the repayment amount and discount rate $(y_{0},d_{0})$ that jointly solve the equations
    ${G(y_{0})+\frac{1-F(x_{0})}{1-F(y_{0})}\frac{d_{0}}{\rho}\frac{1-s(d^{*})\rho}{1-s(d^{*})d^{*}}=1,\quad\frac{1-F(y_{0})}{1-F(x_{0})}\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}=\frac{s(d_{0})+s^{\prime}(d_{0})d_{0}}{\rho s^{\prime}(d_{0})},}$
    which have a unique solution with $d_{0}\in(0,d^{*})$. If the consumer repays $y_{0}$, the lender fixes the repayment amount at $y_{t}=y_{0}$ and the discount rate at $d_{t}=d^{*}$ for all subsequent periods.
  • (b) if $x_{0}\geq\bar{x}$, the optimal policy is to set the repayment amount in each period at $y_{t}=x_{0}$ with a fixed discount rate $d^{*}$.
As in the preceding proof, we omit the dependence on the time period $t$. We define $\beta=\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}$. We have $G(\bar{x})=1-\frac{d^{*}}{\rho\beta}$.
Along the lines of the proof of Theorem 3.6, our proof proceeds by showing that the proposed solution satisfies the Bellman equations for $x<\bar{x}$ and $x\geq\bar{x}$ separately. We evaluate these equations for different states, and the following Lemma addresses the existence and uniqueness of solutions to first order conditions of the form
${G(\Upsilon)+\frac{1-F(X)}{1-F(\Upsilon)}\frac{\delta}{\rho}\frac{1-s(d^{*})\rho}{1-s(d^{*})d^{*}}=1,}$ (51)
and
${\frac{1-F(\Upsilon)}{1-F(X)}\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}=\frac{s(\delta)+s^{\prime}(\delta)\delta}{\rho s^{\prime}(\delta.}}$ (52)
Lemma A.9
Let $X\in(0,\bar{x})$. There exists a unique $(\Upsilon,\delta)$ satisfying equations (51)-(52) for $\delta\in(0,d^{*}]$.
Proof. Define $\eta(\delta)=\delta+\frac{s(\delta)}{s^{\prime}(\delta)}$. With some algebra, we can rewrite equations (51)-(52) with $\Upsilon$ as functions of $\delta$: $\Upsilon=G^{-1}\left(1-\frac{\delta}{\eta(\delta)}\right)$ and $\Upsilon=F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta)\right)$, and it is sufficient to show that these functions intersect at a single $\delta$. Defining $M(\delta)=G^{-1}\left(1-\frac{\delta}{\eta(\delta)}\right)-F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta)\right)$, we thus need to prove that for $\delta\in(0,d^{*}]$, $M(\delta)$ crosses the $\delta$-axis exactly once. Since the function $M(\cdot)$ is differentiable, it is sufficient to show that (i) $\lim_{\delta\to 0}M(\delta)<0$; (ii) $\lim_{\delta\to d^{*}}M(\delta)>0$; and (iii) at any point $\delta\in(0,d^{*}]$ where $M(\delta)=0$, $M^{\prime}(\delta)>0$.
Proof of (i): As $\delta\to 0$, $\lim_{\delta\to 0}M(\delta)=G^{-1}(1)-\lim_{\delta\to 0}F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta)\right)$. Since $G(\cdot)$ is monotone increasing with $\lim_{x\to u}G(x)\to\infty$, there exists a $\hat{y}<u$ such that $G(\hat{y})=1$, or $\hat{y}=G^{-1}(1)$. For the second term, $\lim_{\delta\to 0}F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta)\right)=u$, hence $\lim_{\delta\to 0}M(\delta)=\hat{y}-u<0$.
Proof of (ii): $\lim_{\delta\to d^{*}}M(\delta)=G^{-1}(1-\frac{d^{*}}{\eta(d^{*})})-F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(d^{*})\right)=G^{-1}(1-\frac{d^{*}}{\eta(d^{*})})-x_{0}=\bar{x}-x_{0}>0$. The fact that the first expression is reduced to $\bar{x}$ follows from the definition of $\bar{x}$. As for the second expression, when $\delta=d^{*}$, let $\Upsilon=F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(d^{*})\right)$. This implies $\left(1-F(\Upsilon)\right)\rho\beta=\left(1-F(x_{0})\right)\eta(d^{*})$, and by the definition of $d^{*}$, $\eta(d^{*})=\rho\beta$, hence $\Upsilon=F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(d^{*})\right)=x_{0}$.
Proof of (iii): Let $l(\delta)=\frac{\delta}{\eta(\delta)}$, then
${M^{\prime}(\delta)=-\frac{l^{\prime}(\delta)}{G^{\prime}\left(G^{-1}\left(1-\frac{\delta}{\eta(\delta)}\right)\right)}+\frac{\frac{1-F(x_{0})}{\rho\beta}\eta^{\prime}(\delta)}{f\left(F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta)\right)\right)}.}$ (53)
If $M(\delta)=0$, then $G^{-1}\left(1-\frac{\delta}{\eta(\delta)}\right)=F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta)\right)$ and the corresponding $\Upsilon$ is given by $\Upsilon=G^{-1}\left(1-\frac{\delta}{\eta(\delta)}\right)=F^{-1}\left(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta)\right)$. With that, (53) is reduced to ${M^{\prime}(\delta)=-\frac{l^{\prime}(\delta)}{G^{\prime}(\Upsilon)}+\frac{\frac{1-F(x_{0})}{\rho\beta}\eta^{\prime}(\delta)}{f(\Upsilon)}}$. As $G^{\prime}(\Upsilon)=H(\Upsilon)+H^{\prime}(\Upsilon)\Upsilon$ where $H(\cdot)$ is the monotone increasing hazard function, $G^{\prime}(\Upsilon)>H(\Upsilon)$ and
${M^{\prime}(\delta)>-\frac{l^{\prime}(\delta)}{H(\Upsilon)}+\frac{\frac{1-F(x_{0})}{\rho\beta}\eta^{\prime}(\delta)}{f(\Upsilon)}=\frac{1-F(\Upsilon)}{f(\Upsilon)}\left(\frac{1-F(x_{0})}{1-F(\Upsilon)}\rho\beta\eta^{\prime}(\delta)-l^{\prime}(\delta)\right).}$
As $\Upsilon=F^{-1}(1-\frac{1-F(x_{0})}{\rho\beta}\eta(\delta))$, $\frac{1-F(x_{0})}{1-F(\Upsilon)}\rho\beta=\eta(\delta)$, and $M^{\prime}(\delta)>\frac{1-F(\Upsilon)}{f(\Upsilon)}\left(\eta(\delta)\eta^{\prime}(\delta)-l^{\prime}(\delta)\right)$. With some algebra, $\eta(\delta)\eta^{\prime}(\delta)-l^{\prime}(\delta)=\frac{\eta^{3}(\delta)\eta^{\prime}(\delta)-\eta(\delta)+\delta\eta^{\prime}(\delta)}{\eta^{2}(\delta)}$. Since $\eta^{\prime}(\delta)=2-\frac{s^{\prime\prime}(\delta)}{[s^{\prime}(\delta)]^{2}}>2$, $\eta(\delta)\eta^{\prime}(\delta)-l^{\prime}(\delta)>\frac{2\eta^{3}(\delta)+2\delta-\eta(\delta)}{\eta^{2}(\delta)}=\frac{2\delta^{3}(\xi(\delta)+1)^{3}+\delta(\xi(\delta)-1)\xi^{2}(\delta)}{\eta^{2}(\delta)\xi^{3}(\delta)}$, where $\xi(\delta)$ is the demand elasticity. Since $\xi(\delta)>0$ and $\delta\in(0,1)$, $02\delta^{3}(\xi(\delta)+1)^{3}+\delta(\xi(\delta)-1)\xi^{2}(\delta)>$ 0. Hence, $M(\delta)=0$ implies $M^{\prime}(\delta)>0$. $\square$
Proceeding with the proof of Theorem A.8, we verify the Bellman equation for $x<\bar{x}$ (Case 1) and $x\geq\bar{x}$ (Case 2).
Case 1: $x<\bar{x}$. We define ($y_{x},d_{x})$ as the solution to (51)-(52) with $\Upsilon=y_{x},\delta=d_{x}$ and $X=x$. Lemma A.9 guarantees the existence and uniqueness of the solution. We verify that the following equation holds:
$J(x)|_{x<\bar{x}}=$ $x-s(d_{x})d_{x}y_{x}+\rho s(d_{x})\frac{1-F(y_{x})}{1-F(x)}\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}y_{x}$ (54)
$=$ $\max\{\max_{d,y\in[x,\bar{x})}x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\left(y-s(d_{y})d_{y}y_{y}+\rho s(d_{y})\frac{1-F(y_{y})}{1-F(y)}\beta y_{y}\right),\max_{d,y\in[\bar{x},u)}x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y\},$ (55)
Here, ($y_{y},d_{y})$ is the solution to (51)-(52) with $\Upsilon=y_{y},\delta=d_{y}$ and $X=y$ (where $\Upsilon=y_{y}$ and $\delta=d_{y}$ satisfy the first order conditions for maximizing $\Pi_{y}(\Upsilon,\delta)=\Upsilon-s(\delta)\delta\Upsilon+\rho s(\delta)\frac{1-F(\Upsilon)}{1-F(y)}\beta\Upsilon$, with the current state being $y$). We further define $C(y)=\Pi_{y}(y_{y},d_{y})$, which equals to the term in round brackets in equation (55). The existence and uniqueness of $(y_{y},d_{y})$ follow from Lemma A.9.
Proof. The future continuation value in the Bellman equation depends on the range $y$ is in. We consider two cases: $y\in[x,\bar{x})$ (Case 1.1) and $y\in[\bar{x},u)$ (Case 1.2).
Case 1.1: $y\in[x,\bar{x})$. We prove that
${\max_{d,y\in[x,\bar{x})}x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}C(y)\leq x-s(d_{x})d_{x}y_{x}+\rho s(d_{x})\frac{1-F(y_{x})}{1-F(x)}\beta y_{x}.}$
Consider the optimization problem
${\max_{d,y\in[x,\bar{x})}\Pi_{1}(d,y):=x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}C(y).}$ (56)
For all $y\leq\bar{x}$, $C(y)\leq\beta y_{y}$, as
${C(y)=y-s(d_{y})d_{y}y_{y}+\rho s(d_{y})\frac{1-F(y_{y})}{1-F(y)}\beta y_{y}=y+[\eta(d_{y})-d_{y}]s(d_{y})y_{y}\leq y_{y}+[\eta(d^{*})-d^{*}]s(d^{*})y_{y}=\beta y_{y},}$
where the inequality follows from the fact that the derivative of $\eta(d)-d$ is $01-\frac{s^{\prime\prime}(d)}{[s^{\prime}(d)]^{2}}>$ 0. Since $d_{y}=\eta^{-1}\left(\rho\beta\frac{1-F(y_{y})}{1-F(y)}\right)$ and $d^{*}=\eta^{-1}(\rho\beta)$, $d^{*}>d_{y}$. Hence, for $y\in[x,\bar{x})$,
${\Pi_{1}(y,d)=x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}C(y)<x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y_{y}.}$
Extending the set of feasible $y$ value to $y\in[x,u)$ can only improve the objective function value. Therefore,
${\max_{d,y\in[x,\bar{x})}\ x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y_{y}\leq\max_{d,y\in[x,u)}\ x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y_{y}.}$
As $d_{x}$ and $y_{x}$ are the maximizers in $\max_{d,y\in[x,u)}x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y$, we have
${\max_{d,y\in[x,u)}\ x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y_{y}\leq x-s(d_{x})d_{x}y_{x}+\rho s(d_{x})\frac{1-F(y_{x})}{1-F(x)}\beta y_{x}.}$
Case 1.2: $y\in[\bar{x},u)$. We prove $(d_{x},y_{x})$ jointly maximize of ${\max_{d,y\in[\bar{x},u)}\Pi_{2}(d,y)=x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y}$, and that $y_{x}\geq\bar{x}$. We first optimize over $d$. The optimal $d$, denoted by $d_{x}$, satisfies $\eta(d)=\rho\beta\frac{1-F(y)}{1-F(x)}$. We then optimize over $y$. Let $\Pi_{2}(d,y)=x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y.$ We first prove that at $y=\bar{x}$, $\frac{d\Pi_{2}(d_{x},y)}{dy}>0$. This together with $\lim_{y\to u}\Pi_{2}(d_{x},y)<0$ implies that for $y\in[\bar{x},u)$, the solution that maximizes $\Pi_{2}$ is an interior solution that satisfies the first order condition.
We next prove at $y=\bar{x}$, $\frac{d\Pi_{2}(d_{x},y)}{dy}>0$. This is because
${\frac{d\Pi_{2}(d_{x},y)}{dy}\bigg{|}_{y=\bar{x}}=s(d_{x})\rho\beta\frac{1-F(y)}{1-F(x)}\left(-G(\bar{x})-\frac{d_{x}}{\rho\beta}\frac{1-F(x)}{1-F(\bar{x})}+1\right)=s(d_{x})\rho\beta\frac{1-F(y)}{1-F(x)}\left(\frac{d^{*}}{\rho\beta}-\frac{d_{x}}{\rho\beta}\frac{1-F(x)}{1-F(\bar{x})}\right),}$
where the second equality follows from the definition of $\bar{x}$. Since ${\frac{d_{x}}{\rho\beta}\frac{1-F(x)}{1-F(\bar{x})}=\frac{d_{x}}{\eta(d_{x})},\ \frac{d^{*}}{\rho\beta}=\frac{d^{*}}{\eta(d^{*})}}$,
${\frac{d\Pi_{2}(d_{x},y)}{dy}\bigg{|}_{y=\bar{x}}=s(d_{x})\rho\beta\frac{1-F(y)}{1-F(x)}\left(\frac{d^{*}}{\eta(d^{*})}-\frac{d_{x}}{\eta(d_{x})}\right).}$
For $x<\bar{x}$, $d_{x}<d^{*}$, as ${\rho\beta\frac{1-F(\bar{x})}{1-F(x)}=\eta(d_{x})}$ and ${\rho\beta=\eta(d^{*}).}$ Since $\eta(\cdot)$ is monotone increasing, $d_{x}<d^{*}$. Also, $\frac{d}{\eta(d)}=\frac{1}{1+\frac{1}{\xi(d)}}$ is monotone increasing when $\xi(\cdot)$ (the demand elasticity) is monotone increasing. Hence, $\frac{d\Pi}{dy}|_{\bar{x}}=s(d_{x})\rho\beta\frac{1-F(y)}{1-F(x)}\left(\frac{d^{*}}{\eta(d^{*})}-\frac{d_{x}}{\eta(d_{x})}\right)>0$, while the same derivative is negative as $y\to u$.
[…]
This completes the proof for Case 1 ($x<\bar{x}$). $\square$
Case 2: $x\geq\bar{x}$. We verify that the following equation holds:
${J(x)|_{x\geq\bar{x}}=\frac{1-s(d^{*})d^{*}}{1-s(d^{*})\rho}x=\max_{d,y\in[x,u)}\ x-s(d)dy+\rho s(d)\frac{1-F(y)}{1-F(x)}\beta y.}$
Proof. We first optimize over $d$. The optimal $d$, denoted by $d_{x}$, should satisfy $\rho\beta\frac{1-F(y)}{1-F(x)}=\eta(d)$, so $d_{x}=\eta^{-1}\left(\rho\beta\frac{1-F(y)}{1-F(x)}\right)$. We next optimize over $y$. Let $\Pi_{2}(d_{x},y)=x-s(d_{x})d_{x}y+\rho s(d_{x})\frac{1-F(y)}{1-F(x)}\beta y$, then
${\frac{d\Pi_{2}(d_{x},y)}{dy}=s(d)\rho\beta\frac{1-F(y)}{1-F(x)}\left(-G(y)-\frac{d_{x}}{\rho\beta}\frac{1-F(x)}{1-F(y)}+1\right).}$
Now, $\frac{d_{x}}{\rho\beta}\frac{1-F(x)}{1-F(y)}=\frac{d_{x}}{\eta(d_{x})}=\frac{1}{1+1/\xi(d_{x})}$, where $\xi(d)$ is the demand elasticity. Let $\lambda(x)=\frac{1}{1+1/x}$, $\Lambda(y)=G(y)+\frac{d_{x}}{\rho\beta}\frac{1-F(x)}{1-F(y)}$, $\lambda(\xi(d_{x}))=\frac{d_{x}}{\rho\beta}\frac{1-F(x)}{1-F(y)}$, and $\Lambda(y)=G(y)+\lambda\left(\xi\left(\eta^{-1}\left(\rho\beta\frac{1-F(y)}{1-F(x)}\right)\right)\right)$. With these definitions,
${\frac{d\Pi_{2}(d_{x},y)}{dy}=s(d_{x})\rho\beta\frac{1-F(y)}{1-F(x)}\left(1-\Lambda(y)\right).}$ (57)
By the inverse function theorem,
${\frac{d\ \lambda\left(\xi\left(\eta^{-1}\left(\rho\beta\frac{1-F(y)}{1-F(x)}\right)\right)\right)}{dy}=-\frac{\xi^{\prime}(d_{x})}{(1+\xi(d_{x}))^{2}\eta^{\prime}(d_{x})}\frac{\rho\beta f(y)}{(1-F(x))}.}$
Recall that $H(y)=\frac{f(y)}{1-F(y)}$ is the hazard rate of the income distribution evaluated at $y$. Now,
$\frac{d\Lambda(y)}{dy}$ $=\frac{f(y)}{1-F(y)}+H^{\prime}(y)y-\frac{\xi^{\prime}(d_{x})}{(1+\xi(d_{x}))^{2}\eta^{\prime}(d_{x})}\frac{\rho\beta f(y)}{(1-F(x))}$
$\geq\frac{f(y)}{1-F(y)}-\frac{\xi^{\prime}(d_{x})}{(1+\xi(d_{x}))^{2}\eta^{\prime}(d_{x})}\frac{\rho\beta f(y)}{(1-F(x))}$
$\geq\frac{f(y)}{1-F(y)}\left(1-\frac{\xi^{\prime}(d_{x})\eta(d_{x})}{(1+\xi(d_{x}))^{2}\eta^{\prime}(d_{x})}\right),$
where the first inequality follows from $H^{\prime}(y)\geq 0$ and the second inequality follows from the fact that $x\leq y$, hence $1-F(x)\geq 1-F(y)$.
To complete the proof, we show that for all $d\in(0,1)$, $\frac{d\Lambda(y)}{dy}\geq 0$. For all $d\in(0,1)$, $\xi^{\prime}(d)\eta(d)>0$, $(1+\xi(d))^{2}\eta^{\prime}(d)>0$, and $s^{\prime}(d)>0$, $s^{\prime\prime}(d)<0$. Thus,
$(1+\xi(d))^{2}\eta^{\prime}(d)-\xi^{\prime}(d)\eta(d)=\left(1+\frac{ds^{\prime}(d)}{s(d)}\right)^{2}\left(2-\frac{s^{\prime\prime}(d)}{(s^{\prime}(d))^{2}}\right)-\left(d+\frac{s(d)}{s^{\prime}(d)}\right)\frac{\left(s(d)(s^{\prime}(d)+s^{\prime\prime}(d)d)-d(s^{\prime}(d))^{2}\right)}{(s(d))^{2}}$
$=\frac{s(d)+ds^{\prime}(d)}{(s(d)s^{\prime}(d))^{2}}\left((s^{\prime}(d))^{2}(s(d)+3ds^{\prime}(d))-s^{\prime\prime}(d)(s(d)+ds^{\prime}(d)(1+s(d)))\right)>0,$
which implies $\frac{\xi^{\prime}(d_{x})\eta(d_{x})}{(1+\xi(d_{x}))^{2}\eta^{\prime}(d_{x})}<1$.
It follows that $\frac{d\Lambda(y)}{dy}>0$, so $\Lambda(y)$ is monotone increasing and at $y=x\geq\bar{x}$,
${\frac{d\Pi_{2}(d_{x},y)}{dy}\bigg{|}_{y=x}=s(d^{*})\rho\beta\left(-G(x)+\frac{d^{*}}{\rho\beta}+1\right)\leq 0.}$
To complete the proof, we show that for $y>x$, $\frac{d\Pi_{2}}{dy}$, given by equation (57), is negative, which implies that the maximum is achieved at $y=x$. This follows from the fact that as shown above, $\Lambda(y)$ is strictly increasing so $(1-\Lambda(y))<0$ for $y>x$ while the other terms in (57) are positive. We thus conclude that when $x\geq\bar{x}$, $(x,d^{*})$ maximizes $\Pi_{2}(d,y)$, which completes the proof of Case 2. $\square$
A.12 Proof of Proposition 3.7
We compare the dynamic and hybrid architectures separately for $x_{0}\in(0,\bar{x})$ and $x_{0}\in[\bar{x},u)$. Table 1 compares the loan amounts at $t=0$ and the retention probabilities from $t=0$ to $t=1$.
Loan Amount ($t=0$) Retention Probability from $t=0$ to $t=1$
Dynamic, $x_{0}\in(0,\bar{x})$ $(1-F(\bar{x}))d^{*}\bar{x}$ $s((1-F(\bar{x}))d^{*})\dfrac{1-F(\bar{x})}{1-F(x_{0})}$
Hybrid, $x_{0}\in(0,\bar{x})$ $\dfrac{1-F(\bar{x})}{1-F(x_{0})}d^{*}\bar{x}$ $s\left(\dfrac{1-F(\bar{x})}{1-F(x_{0})}d^{*}\right)\dfrac{1-F(\bar{x})}{1-F(x_{0})}$
Dynamic, $x_{0}\in[\bar{x},u)$ $(1-F(\bar{x}))d^{*}\bar{x}$ $s((1-F(\bar{x}))d^{*})$
Hybrid, $x_{0}\in[\bar{x},u)$ $d^{*}x_{0}$ $s(d^{*})$
Table 1: Comparison of loan amounts at $t=0$ and transition probabilities to $t=1$ for the dynamic and hybrid architectures
First, the coverage at $t=0$ is the same under both architectures as the consumer always gets a loan. Further, for both $x\in(0,\bar{x})$ and $x\in[\bar{x},u)$, the loan amount is clearly larger under the hybrid architecture.
The retention probability from $t=0$ to $t=1$ is the probability of the consumer accepting and repaying the loan. When $x_{0}\in(0,\bar{x})$, the repayment amount is $\bar{x}$ under both architectures, so the probability of successful repayment is the same, however the discount rate, hence the retention probability, is higher under the hybrid policy.
For $x\in[\bar{x},u)$, the repayment amount, hence the repayment probability, is again the same under both architectures. However, the discount rate under the hybrid architecture, $d^{*}$, is higher than the dynamic policy’s $(1-F(\bar{x}))d^{*}$, again resulting in a higher retention probability.
From period $t=1$ onward, both architectures stabilize with the same retention probability $s(d^{*})$. Regarding the loan amounts, for $x_{0}\in(0,\bar{x})$, they are the same under both architectures. For $x_{0}\in[\bar{x},1)$, the hybrid architecture yields a loan amount of $d^{*}x_{0}$, higher than the dynamic policy’s $d^{*}\bar{x}$. $\square$
A.13 Proof of Proposition 4.1
We start by proving a general property of the lender’s expected NPV, $\operatorname{\mathbb{E}}(\Pi(\lambda))=\frac{(d^{*})^{\alpha+1}}{\alpha}[1-F_{\lambda}(\bar{x}(\lambda))]^{\alpha+1}\bar{x}(\lambda)$, when the income distribution $F_{\lambda}(\cdot)$ and the NPV are parameterized by an arbitrary parameter $\lambda$: if for all $x$, $1-F_{\lambda}(x)$ is quasi-convex in $\lambda$, then $\operatorname{\mathbb{E}}(\Pi(\lambda))$ is quasi-convex in $\lambda$. In particular, this is satisfied by the Beta distribution.
To prove this property, first note that quasi-convexity of $1-F_{\lambda}(x)$ means that for all $\lambda_{1}<\lambda_{2}$ and $\gamma\in[0,1]$, if we define $\lambda=\gamma\lambda_{1}+(1-\gamma)\lambda_{2}$, then for all $x$, $1-F_{\lambda}(x)\leq\max\{1-F_{\lambda_{1}}(x),1-F_{\lambda_{2}}(x)\}$.
If $1-F_{\lambda_{1}}(x)\geq 1-F_{\lambda_{2}}(x)$, then $1-F_{\lambda}(x)\leq 1-F_{\lambda_{1}}(x)$. Moreover,
${\operatorname{\mathbb{E}}(\Pi(\lambda))=\frac{(d^{*})^{\alpha+1}}{\alpha}[1-F_{\lambda}(\bar{x}(\lambda))]^{\alpha+1}\bar{x}(\lambda)\leq\frac{(d^{*})^{\alpha+1}}{\alpha}[1-F_{\lambda_{1}}(\bar{x}(\lambda))]^{\alpha+1}\bar{x}(\lambda)\leq\frac{(d^{*})^{\alpha+1}}{\alpha}[1-F_{\lambda_{1}}(\bar{x}(\lambda_{1}))]^{\alpha+1}\bar{x}(\lambda_{1})=\operatorname{\mathbb{E}}(\Pi(\lambda_{1})),}$
where the second inequality follows from $\bar{x}(\lambda_{1})$ being the maximizer of the profit function under distribution $F_{\lambda_{1}}$. Hence, $\operatorname{\mathbb{E}}(\Pi(\lambda))\leq\operatorname{\mathbb{E}}(\Pi(\lambda_{1}))$. Similarly, if $1-F_{\lambda_{1}}(x)<1-F_{\lambda_{2}}(x)$, by quasi convexity, $1-F_{\lambda}(x)\leq 1-F_{\lambda_{2}}(x)$, and following the same argument, $\operatorname{\mathbb{E}}(\Pi(\lambda))\leq\operatorname{\mathbb{E}}(\Pi(\lambda_{2}))$. Thus, if $1-F_{\lambda}(x)\leq\max\{1-F_{\lambda_{1}}(x),1-F_{\lambda_{2}}(x)\}$, then $\operatorname{\mathbb{E}}(\Pi(\lambda))\leq\max\{\operatorname{\mathbb{E}}(\Pi(\lambda_{1})),\operatorname{\mathbb{E}}(\Pi(\lambda_{2}))\}$. It follows that quasi-convexity of $1-F_{\lambda}(x)$ for all $x$ implies quasi-convexity of $E(\Pi(\lambda))$.
For the Beta distribution with $\lambda=a=b$ (i.e., a fixed mean), let $\lambda=\gamma\lambda_{1}+(1-\gamma)\lambda_{2}$ for $\gamma\in[0,1]$. Then, $F_{\lambda}(x)\geq\min\{F_{\lambda_{1}}(x),F_{\lambda_{2}}(x)\}$ because decreasing $\lambda$ is a mean-preserving spread (Rothschild and Stiglitz [1978]). Hence, $1-F_{\lambda}(x)\leq\max\{1-F_{\lambda_{1}}(x),1-F_{\lambda_{2}}(x)\}$, and as shown above, the expected NPV is quasi-convex in $\lambda$.
We are now ready to show that the NPV is U-shaped. As $\lambda\to 0$, the distribution becomes a standard Bernoulli distribution and $\operatorname{\mathbb{E}}(\Pi(\lambda=0))=\frac{2^{-1-\alpha}}{\alpha}{\left(\frac{{\alpha(-1+d^{1+\alpha})\rho}}{{(1+\alpha)(-1+d^{\alpha}\rho)}}\right)^{1+\alpha}}$. As $\lambda\to\infty$, the distribution becomes degenerate with $P(\theta=0.5)=1$ and $\lim_{\lambda\to\infty}\operatorname{\mathbb{E}}(\Pi(\lambda))=\frac{1-(d^{*})^{\alpha+1}}{1-(d^{*})^{\alpha}\rho}\frac{1}{2}$. When $\lambda=1$, the distribution is uniform and $\operatorname{\mathbb{E}}(\Pi(\lambda=1))=\frac{(d^{*})^{\alpha+1}}{\alpha}(1+\alpha)^{1+\alpha}(2+\alpha)^{-(2+\alpha)}<\min\{\operatorname{\mathbb{E}}(\Pi(\lambda=0)),\lim_{\lambda\to\infty}\operatorname{\mathbb{E}}(\Pi(\lambda))$ for $\alpha\in(0,1)$. Since $\operatorname{\mathbb{E}}(\Pi(\lambda))$ is continuous and quasi-convex, it follows that $\operatorname{\mathbb{E}}(\Pi(\lambda))$ first decreases and then increases in $\lambda$ (where $\lambda=a=b$) or, equivalently, it first decreases and then increases in the income variance $\frac{1}{4(2a+1)}$. $\square$